### v1

# Gibbs states and Brownian models for coexisting haze and cloud droplets

Abstract
Clouds microphysics describes the formation and evolution of cloud droplets, rain, and ice particles. It is among the most critical factors in determining the cloud’s size, lifetime, precipitation, and radiative effect. Among all cloud types, the small clouds, characterized by weak updrafts, that are close to the haze-to-cloud transition pose challenges in measuring them and understanding their properties. They are superabundant but hard to capture by satellites and often falsely regarded as aerosols. Köhler’s theory explains droplet activation and their growth by condensation at the earliest stages of cloud development. It fully describes the thermodynamic state of a single drop but falls short when explaining the collective behavior of large populations of particles. This is especially important when the supersaturation pool is limited. We present an analytical framework to extend Köhler’s theory to coexisting cloud droplets. Our results suggest hysteresis and asymmetry in the process of droplet activation and deactivation. The turbulent nature of clouds is incorporated into our model formulation as Brownian noise to provide explicit droplet size distributions and activation timescales. The theoretical findings are confronted with experimental data stemming from laboratory clouds created in the Pi convection chamber, suggesting a new way of understanding haze-to-cloud transitions and small cloud formation processes.
INTRODUCTION
Clouds play a key role in the climate energy balances and in the water cycle Bengtsson (2010). A large part of the uncertainties in climate predictions is attributed to limitations in our understanding of cloud physics and therefore, how to parameterize such clouds in global climate models Zelinka et al. (2020). Out of all cloud types, a great part of the cloud research challenges is attributed to small, warm clouds that are highly abounded but hard to measure. Many of these clouds are too small relative to earth-observing satellites’ resolution, and/or their optical signature is too weak Koren et al. (2007); Várnai and Marshak (2009); Eytan et al. (2020).
The cloud’s droplet size distribution (DSD) is among the most important microphysical properties. The DSD properties are coupled to all dynamical, turbulent, optical, and stochastic cloud processes. DSDs affect the cloud’s evolution in time, rain processes, cloud depth, size, and lifetime Khain and Pinsky (2018). When supersaturation is high enough, cloud condensation nuclei (CCNs) are activated into cloud droplets and grow by condensation. This is the main mechanism for droplet growth at the early stages of a warm cloud. If the cloud’s dynamics is weak (shallow clouds) or if the aerosol concentration is high, growth droplets by condensation could be the main mechanism throughout the whole cloud’s lifetime as droplets do not become large and varied enough to support efficient autoconversion of droplets by the stochastic processes of collision and coalescence (Rogers and Yau, 1989, §7-8).
In warm clouds that are controlled by weak updrafts and low supersaturations, only a small part of the aerosol (the large and most hygroscopic) are activated. A competition between the droplet curvature (Kelvin) effect, which suppresses activation, and the solute (Raoult) effect, which reduces the critical supersaturation for activation, Köhler (1936) allows for suspended particles in air parcels to coexist in two main thermodynamical states known as: haze and activated droplet Khain and Pinsky (2018). While hazy particles are in stable equilibrium, activated particles harvest the available water vapor and grow until equilibrium is restored.
The mechanisms of activation and deactivation are, however, not symmetric. In order to activate a cloud droplet, it is necessary to attain a critical level of supersaturation $\lambda_{K}$, typically well above $0\%$. When supersaturation is decreased to $\lambda_{K}$ in the deactivation process, an activated cloud droplet is big enough to still experience diffusion of water vapor onto its surface and, hence, it is expected that this positive feedback makes the threshold for deactivation strictly lower than that of $\lambda_{K}$. Consequently, activation-deactivation undergoes hysteresis, which has been previously reported in the cloud physics literature Korolev and Mazin (2003); Arabas and Shima (2017). From a theoretical perspective, in the work of Arabas and Shima (2017) it was shown on analytical grounds that the activation process of a cloud particle undergoes a saddle-node bifurcation and, when coupled to supersaturation evolution, hysteresis is numerically observed in a parcel model and attributed to kinetic limitations and the presence of a cusp bifurcation (Kuznetsov, 2004, §8.2).
Another level of complexity in the determination of microphysical properties concerns clouds living in eminently turbulent environments. Indeed, turbulence heavily influences the life cycle of clouds; see Shaw (2003) for a review on the topic. In particular, the incorporation of turbulent models provides the means to explain droplet size spectra broadening Sardina et al. (2015); Grabowski and Abade (2017), as well as describe the mixing of the cumulus clouds with their exterior Abade, Grabowski, and Pawlowska (2018); Eytan et al. (2020) and parameterize mixed-phase clouds in global circulation models Field et al. (2014); Furtado et al. (2016). In the context of warm clouds, localized temperature gradients and turbulence at the smallest scales provoke supersaturation fluctuations that allow activation of haze particles even in mean-subsaturated environments Prabhakaran et al. (2020), and similarly, cloud particles susceptible of deactivating Abade, Grabowski, and Pawlowska (2018). Thus, microphysical fluctuations are a key feature in obtaining faithful descriptions of droplet size spectra.
This research explores the activation and deactivation of cloud droplets in warm clouds, offering new theoretical perspectives. We aim to find analytical formulas for describing the size distribution of droplets (DSDs) when both small cloud droplets and haze particles coexist, assuming condensation as the only growth mechanism. Similar to previous work Arabas and Shima (2017), we analyze how droplet growth due to condensation changes dramatically as the surrounding air’s supersaturation level varies. To the Köhler instability we add a sink term in growth rate which aims at (i) modeling the supersaturation budget, and (ii) parameterising other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al. (2014); Altaratz et al. (2021). The resulting equation describing condensation growth for droplets of a uniform size allows for multiple stable states to co-exist besides the ”haze” state, meaning the system exhibits multistability. Although the haze state remains stable, this multistability suggests that activated droplets can get stuck around a specific size due to the balancing act between two factors: the Köhler instability and mechanisms that decrease supersaturation or remove particles; see also Korolev and Mazin (2003); Arabas and Shima (2017); Chandrakar et al. (2020).
However, turbulence effects throw a wrench in the picture. The stable states become “fuzzy states” that are encountered statistically (metastable) at various rates of occurrence. This means that particles can switch between activated and deactivated states, similar to how Brownian particles move around in a bumpy landscape. These turbulent jolts also significantly impact the activation-deactivation process into the form of various possible hysteresis loops (see Fig. 5 below). As a result, droplets can become activated at lower levels of supersaturation compared to the standard Köhler threshold.
Finally, our framework allows us to derive analytical formulas for the size distribution of droplets (DSDs) within a cloud volume, accounting for these stochastic turbulent effects. In particular, our stochastic model predicts that droplet size distributions in Pi cloud convection chamber experiments behave like Gibbs states. It is shown that this prediction closely matches the actual DSDs observed in these experiments (see Fig. 6 below).
RESULTS
Köhler theory
The theory of Köhler is the thermodynamical backbone to explain the formation of cloud droplets out of suspended aerosols in the atmosphere. It establishes the critical relative humidity threshold beyond which suspended particles activate and grow by diffusion (condensation) Köhler (1936). The critical threshold is above saturation (hence, the supersaturation term is used), and it depends on the initial size and chemical composition of the particle in question. These particles are, according to Köhler, present in two distinct/dichotomical thermodynamic states: The first is known as haze and it is observed when the aerosol is humidified, although remains at equilibrium with its moist environment. The second state is referred to as activated droplet, whereby the particle harvests the available moisture by diffusion and, in principle, grows indefinitely. However, in nature, the available water molecules that are in supersaturation are limited, and in many cases, droplets and haze coexist in the cloud, and particles can oscillate between the two states. Particularly, in cases of small, warm clouds Hirsch et al. (2017); Altaratz et al. (2021).
The critical relative humidity threshold needed for activation of a given aerosol with a solubility constant $k$ and a dry radius $r_{d}$ is analytically obtained using the Köhler’s equation Köhler (1936); Rogers and Yau (1989). Here, to simplify some derivations, $\lambda$ denotes the part of the relative humidity above $100\%$. If $r(t)$ denotes the radius of a given droplet at time $t$, its evolution is approximated by Köhler (1936); Arabas and Shima (2017):
$\begin{cases}\dot{r}=\frac{D}{r}(\lambda-\lambda_{eq}(r))\\ \lambda_{eq}(r)=\frac{A}{r}-\frac{B}{r^{3}},\end{cases}$ (1)
where $D$ is the diffusivity constant, $A$ relates to the water surface tension, $B=kr_{d}^{3}$, $\lambda$ is the ambient supersaturation and $\lambda_{eq}(r)$ determines the equilibrium supersaturation at the surface of a droplet of radius $r$; commonly known as Köhler curve Wallace and Hobbs (2006). Equation (1) shows that the particle is in equilibrium when $\lambda=\lambda_{eq}(r)$. To simplify the proposed model, we replace $r$ with the variable $X=r^{2}/2D$, which has time-units and could be viewed as relaxation time for diffusion. With this change of variable, Eq. (1) translates into the nondimensional form:
$\begin{cases}\dot{X}=\lambda-f(X),\\ f(X)=A(2DX)^{-1/2}-B(2DX)^{-3/2},\end{cases}$ (2)
with $f(X)$ denoting the equilibrium supersaturation $\lambda_{eq}(r)$ as function of $X$. A rudimentary analysis of $f(X)$ reveals a single global maximum located at $X_{K}=3B/(2DA)$, and with critical supersaturation value $\lambda_{K}=f(X_{K})=(4A^{3}/27B)^{1/2}$, see also (Wallace and Hobbs, 2006, §6.1). An example of Köhler curve is shown in Fig. 1a) whose parameters are given in Table 1. This curve reaches its maximum at $X_{K}=6\cdot 10^{-3}\mathrm{s}$, or at $r^{2}=0.48\mathrm{\mu m}^{2}$.
Table 1: Köhler curve’s parameters for NaCl.
$k$ $r_{d}$ $(\mu\mathrm{m})$ $A$ $(\mu\mathrm{m})$ $D$ ($\mu\mathrm{m}^{2}s^{-1}$)
$\mathrm{NaCl}$ $1.28$ $5\times 10^{-2}$ $10^{-3}$ $40$
For a given dry radius $r_{d}$, when the supersaturation $\lambda$ exceeds the corresponding $\lambda_{K}$, the particle is activated into a cloud droplet and grows by diffusion. On the contrary, for a given ambient supersaturation $\lambda$, particles whose radii are less than $(2DX_{u})^{1/2}$ remain stable in a haze state; see Fig. 1a).
While Köhler theory is established for single particles, clouds consist of large families of particles coexisting in the same humidity environment. When a collection of droplets consumes the available supersaturation, the particles may cease growing by condensation, and the droplet size distributions (DSDs) stagnate and stop being displaced towards larger sizes. Here we modify the Köhler equation to the case of many particles in conditions of small warm clouds, when the available supersaturation is a limit factor. The proposed model extends the condensational growth equation—governed by the Köhler curve— by coupling all particles to supersaturation consumption and by adding stochastic disturbances, to account for (micro) small-scale turbulent effects.
Figure 1: Panel a): Köhler curve associated with $\mathrm{NaCl}$; see Table 1 for the parameter values. A dry radius of $50\mathrm{nm}$ is prescribed. $\lambda_{K}$ denotes Köhler’s critical supersaturation and $\lambda$, an example of supersaturation. Panel b): In black, the Köhler curve associated with $\mathrm{NaCl}$ as Panel a). The cyan to blue curves show different M-Köhler curves associated with Eq. (6) for different values the parameter $\beta(N)$ as indicated in the legend. For the cyan curve are shown furthermore its intersection with a given supersaturation $\lambda$ in between $\lambda_{c}$ and $\lambda_{h}$. Each filled dot corresponds to a stable equilibrium while the the empty dot corresponds to the unstable equilibrium. Panel c): The potential function associated with the cyan M-Köhler curve for the supersaturation value as indicated in the title. The solid/empty dots locate the stable/unstable equilibria.
Multistable Köhler curves
In cases with relatively small supersaturation, as particles grow, they may consume the available supersaturation and therefore reduce their growth rate until reaching stagnation. In the realm of single-particle models these effects can be modeled by the introduction of a sink function into Eq. (1):
[…]
in which $g$ denotes the sink term. The function $f-g$ is called a M-Köhler curve hereafter. The term $g$ should remain negative and should decrease as $X$ increases. The sink term in growth rate aims at (i) modeling the supersaturation budget, and (ii) parameterizing other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al. (2014); Altaratz et al. (2021).
The sink term can also relate to first-order fluctuations in supersaturation. Indeed, when a sizable amount of particles coexist in the same cloud parcel, the constant supersaturation assumption has to be amended taking account changes in $\lambda$ in the course of time, assuming $\lambda$ in Eq. (3) to be time-dependent. At a first order, we are concerned with describing the evolution of $\lambda(t)=\lambda+\lambda^{\prime}(t)$ with $\lambda^{\prime}(0)=0$, i.e. we are seeking for a model of the rate of change of the fluctuations $\lambda^{\prime}$ around a given value $\lambda$.
Following Arabas and Shima (2017), the rate of change $\lambda^{\prime}$ can be approximated by the ratio $-\dot{\rho_{v}}/\rho_{vs}$, in which $\rho_{v}$ denotes the ambient vapour density away from the droplet while, $\rho_{vs}$, denotes the vapour density at saturation. Then, by expressing $\rho_{v}$ as $\rho_{v}=4\pi N\rho_{w}r^{3}/3$, accounting for the density of liquid water $\rho_{w}$ and the droplet volume concentration $N$ of identical (monodisperse) particles in the cloud parcel, we arrive at:
$\dot{\lambda^{\prime}}=-\beta(N)\dot{\left(X^{\frac{3}{2}}\right)},$ (4)
when written in the variable $X=r^{2}/(2D)$, where the constant $\beta(N)$ is given by
$\beta(N)=\frac{2^{7/2}D^{3/2}\pi\rho_{w}}{3\rho_{vs}}N;$ (5)
see (Arabas and Shima, 2017, Eqns. (9) and (10)).
The integration of Eq. (4) yields to $\lambda(t)=\lambda-\beta(N)X^{3/2}$ which eventually leads to the following condensational growth model:
$\dot{X}=\lambda-f(X)-\beta(N)X^{3/2}.$ (6)
Thus, the term $-\beta(N)X^{3/2}$ acts as the sink function $g$ in Eq. (3): it is negative and decreases as $X$ increases. For large values of $X$, it saturates the particle growth. A few M-Köhler curves are shown in Fig. 1b), for different $\beta(N)$-values.
A careful examination of Eq. (6) reveals that Köhler’s critical supersaturation $\lambda_{K}$ and radius are modified due to the sink function. In fact, the function $f(X)+\beta(N)X^{3/2}$ can now admit a local minimum additional to the maximum exhibited by the Köhler curve $f$. These are obtained as zeros of the following depressed cubic polynomial equation:
$-3\beta(N)X^{3}+A(2D)^{-1/2}X-3B(2D)^{-3/2}=0.$ (7)
Sufficient conditions for Eq. (7) to have two positive solutions are easily derivable. Actually, a simple proposition inferring the existence of two positive steady states $X^{\ast}_{h}$ and $X^{\ast}_{c}$ (with $X^{\ast}_{c}\geq X^{\ast}_{h}$) to condensational growth model analogues to Eq. (6) for a sink function $-\beta(N)X^{\alpha}$ ($\alpha>1$) is derived in in Methods (see Proposition 1 therein). In any event, as shown in Fig. 1b) the locations of $X^{\ast}_{h}$ and $X^{\ast}_{c}$ depend on the strength of the sink function’s coefficient $\beta(N)$.
We define then the following supersaturation values $\lambda_{h}=f(X^{\ast}_{h})-g(X^{\ast}_{h})$, which corresponds to the local maximum of $f-g$, and $\lambda_{c}=f(X^{\ast}_{c})-g(X^{\ast}_{c})$, corresponding to its local minimum. In particular, $\lambda_{h}\geq\lambda_{c}$.
Due to the presence of this new local minimum (compared to the Köhler curve), as supersaturation $\lambda$ varies, the number of equilibria varies through two saddle-node bifurcations, as sumarized in Fig. 2.
Figure 2: Going through multiple equilibria as $\lambda$ varies. Two saddle-node (SN) bifurcations occur, one at $\lambda=\lambda_{c}$ and one $\lambda=\lambda_{h}$. For $\lambda<\lambda_{c}$, there is only a hazy stable equilibrium $X_{h}$ (blue dot in the leftmost upper diagram). As one crosses from below the SN bifurcation point at $\lambda=\lambda_{c}$, two other equilibria emerge for a total of 3 equilibria coexisting for $\lambda_{c}<\lambda<\lambda_{h}$ (middle upper diagram): an activated stable equilibrium $X_{c}$ (red dot), an unstable equilibrium $X_{u}$ (open circle), and a haze equilibrium (blue dot). At the SN bifurcation occurring at $\lambda=\lambda_{h}$, $X_{u}$ and $X_{h}$ collide (purple dot on the rightmost lower diagram), and only the activated equilibrium survives as $\lambda>\lambda_{h}$ (red dot in the rightmost upper diagram).
These two saddle-node bifurcations provide the nonlinear backbone for hysteresis to take place as discussed in Hysteresis effects, below. These hysteresis effects are well-known to be revealed in presence of stochastic disturbances; see e.g. Berglund (2011). We discuss below the physical motivation of introducing such disturbances in Eq. (6) or more generally, Eq. (3).
Brownian models and turbulent effects
In a developing cloud, turbulent updrafts and mixing with its neighboring air provoke supersaturation fluctuations that naturally affect droplet growth and, consequently, the collective droplet size spectra Grabowski and Abade (2017); Abade, Grabowski, and Pawlowska (2018). Stochastic models for the velocity fluctuations and the supersaturation fields have been considered in previous studies and have shown various degrees of relevance Sardina et al. (2015); Siewert, Bec, and Krstulovic (2017). For a cloud parcel subject to turbulent updrafts, the authors in Sardina et al. (2015) prove that activated droplet size variance increases, at short times, proportionally to the square-root of time, analogous to Einstein’s diffusion formula. Their analytical findings are contrasted with direct numerical simulations of turbulence obeying the incompressible Navier-Stokes equation. In spite of this diffusion result with diverging variance, in Siewert, Bec, and Krstulovic (2017), the authors find a statistical steady state for droplet size, with DSDs characterized by exponential tails; see their Eq. (4.3). Noteworthy is that to achieve convergence to a steady state, a substantial proportion of the droplets must reach evaporation, possibly if, on average, they experience negative supersaturations.
Supersaturation being the driving parameter, we assume that it evolves around a given value $\lambda$ with fluctuations encoded by a stochastic term possibly depending on the particle size. For the sake of generality, we aim at modeling the inhomogeneous effects of supersaturation fluctuations on droplets of different size. Indeed, this size-dependent fluctuations appear, for instance, when taking the adiabatic limit of fast supersaturation fluctuations, as demonstrated in (Siewert, Bec, and Krstulovic, 2017, Eq. (4.3)). Hence, incorporating different supersaturation sources and sinks alter the expression for the noise law affecting droplet growth. Such noise inhomogeneity is here motivated from the fact that small fluctuations in droplet radius incur greater fluctuations in supersaturation when the particles are bigger. Contrary to haze particles, which are at a stable thermodynamic equilibrium and variations in their radius provoke smaller alterations in the supersaturation budget.
These considerations invite us to parameterize the fluctuations of $\lambda$ by means of a size-dependent stochastic term, $\sigma(X)\dot{W}_{t}$, leading to the following stochastic model:
$\dot{X}=\underbrace{\lambda-f(X)+g(X)}_{\textstyle{-V_{\lambda}^{\prime}(X)}}+\sigma(X)\dot{W}_{t},$ (8)
where $W_{t}$ is a Brownian motion (and $\dot{W}_{t}$ is Gaussian and white).
The function $V_{\lambda}$ is the potential function that collects the nonlinear, deterministic effects in the model. Thus, the solutions $X(t)$ to Eq. (8) can be regarded as Brownian particles embedded in the potential $V_{\lambda}$ defined in Eq. (8). The equilibria of Eq. (3) correspond exactly to the local critical points of the potential function—see Fig. 1c)—and their stability is determined by the local curvature of $V_{\lambda}$ at those points, see e.g. Berglund (2011); Pavliotis (2014). Note that Eq. (8) can always be recast into a form in which the stochastic disturbance term becomes $X$-independent, at the expense of changing the potential function via the Lamperti transformation (Pavliotis, 2014, §3.6). This trick is recalled in Methods as it is also part of our model’s analysis below.
For a family of monodisperse droplets in the same cloud volume, each one experiences different supersaturation fluctuations, albeit with the same mean and variance, as encapsulated by the term $\sigma(X)\dot{W}_{t}$ in Eq. (8). When each supersaturation’s stochastic realization is averaged over a large number of particles, the Fokker-Planck Equation (FPE) associated with Eq. (8) provides the DSD at a given time $t$; see e.g. Pavliotis (2014); Risken (1989). The stationary solution to the FPE is then given analytically and known as the Gibbs state (Pavliotis, 2014, Eq. (4.35)):
$\rho_{\lambda}(X)=Z_{\lambda}^{-1}\frac{\exp\bigg{(}2\int^{X}\frac{\lambda-f(x)+g(x)}{\sigma^{2}(x)}\,\mathrm{d}x\bigg{)}}{\sigma^{2}(X)},$ (9)
where $Z_{\lambda}$ is a normalizing constant. This distribution is a generalization of what is found in (McGraw and Liu, 2006, Eq. (6)). Such equation yields a Weibull distribution for droplet radius $r$, without considering curvature and chemistry effects, i.e. Köhler theory, sink functions or state-dependent noise. One can, thus, derive the Weibull distribution $\rho_{W}$ for the radius $r$ by setting $f=g\equiv 0$ and $\sigma$ constant, where the resulting DSD is integrable only if there is mean subsaturation, $\lambda<0$:
$\rho_{W}(r)=Z_{\lambda}^{-1}re^{\gamma r^{2}},$ (10)
where $\gamma$ is a parameter proportional to $\lambda$, droplet number concentration and liquid water fraction; see (McGraw and Liu, 2006, Eq. (8)). In comparison, $\rho_{\lambda}$ in Eq. (9) can be integrable if $\lambda$ is positive, provided that the sink function $g$ has suitable decay properties detailed in Proposition 2 found in The confining potential in Methods. For instance considering sedimentation, or any size-independent removal mechanism, would provide an integrable DSD; see Krueger (2020) and (Chandrakar et al., 2020, Sections 2.1, 2.2 and 2.3), where the resulting DSDs are an explicit evaluation of $\rho_{\lambda}$ in Eq. (9).
For a broad class of $f$, $g$ and $\sigma$, the Gibbs state $\rho_{\lambda}$ exhibits multimodality intimately related to shape of the potential $V_{\lambda}$. More precisely, the Gibbs state’s modes, the local maxima of $\rho_{\lambda}$, are found by solving the following equation:
$-V^{\prime}_{\lambda}(X)=\frac{1}{2}\sigma^{\prime}(X)\sigma(X).$ (11)
See Noise-induced metastability in Methods for the derivation of this equation. This fundamental equation teaches us how the Gibbs state modes result from the interaction between the noise term and the potential function. The Gibbs state’s modes correspond to metastable states for which the random fluctuation term, $\sigma(X)\dot{W}_{t}$, in Eq. (8) acts as a source for the particles to experience transitions from a statistically typical droplet size to another one; see schematic of Fig. 3.
When the noise is state-independent, i.e. when $\sigma(X)=\sqrt{2\varepsilon}$, then $\sigma^{\prime}(X)=0$ and the metastable states are directly determined by the zeros of $-V_{\lambda}^{\prime}$. In this case, for Eq. (6), the Gibbs states are bimodal when $\lambda_{c}<\lambda<\lambda_{h}$, and unimodal when there is just one underlying stable equilibrium, i.e. for either $\lambda<\lambda_{c}$ or $\lambda>\lambda_{h}$; see Fig. 2. When $\lambda_{c}<\lambda<\lambda_{h}$, there is a non-vanishing probability of encountering haze or activated droplets; see Fig. 3. This property is consistent with realistic situations in which turbulent supersaturation fluctuations make haze susceptible of activating even if on average the Köhler’s threshold is not attained Prabhakaran et al. (2020).
Figure 3: Cartoon of haze-to-droplet stochastic transitions. Here $\lambda$ is kept fixed and the situation depicted corresponds to $\lambda$ in $(\lambda_{c},\lambda_{h})$ where coexist two stable states (haze particle and activated droplet); see Fig. 2.
Our modeling allows for other useful analytic insights. For instance, the expected residence times at the haze or activated state can be precisely determined. They are indeed known to relate to the potential barrier, $V_{\lambda}(X_{u})-V_{\lambda}(X_{h})$ (see Fig. 1c), according to the Kramers’ time formulas Kramers (1940); Berglund (2011). The mean activation time $\tau_{act}$ is then given by (Matkowsky and Schuss, 1981, Eq. (2.19))
$\tau_{act}(\lambda)=\frac{2\pi}{\sqrt{|V_{\lambda}^{\prime\prime}(X_{u})|V_{\lambda}^{\prime\prime}(X_{h})}}e^{\frac{V_{\lambda}(X_{u})-V_{\lambda}(X_{h})}{\varepsilon}}.$ (12)
Therefore, the closer $\lambda$ is to the critical value $\lambda_{h}$ from below, the closer $X_{u}$ is to a collision with $X_{h}$ (see Fig. 2) and the shallower the potential barrier becomes, resulting into more common particle activation. Similarly, deactivation of activated droplets can take place due to fluctuations in supersaturation with a mean deactivation time obtained by (12) in which $X_{h}$ is replaced by $X_{c}$.
Hysteresis effects
Figure 4: Rationale of an hysteresis path. As $\lambda$ varies, the potential landscape changes and the stochastic disturbances can help trigger the critical transition as one approaches the tipping points, here the saddle-node bifurcation points.
Figure 5: Panels a), b) and c): The thin gray line indicates the hysteresis loops for different rates of change of supersaturation $\dot{\lambda}$ as indicated in the title. The noise variance is constant for these three panels as indicated in the legend of Panel a). Panels d), e) and f): The thin gray lines indicate the hysteresis loops for different noise variances as indicated in the titles. The rate of change of supersaturation $\dot{\lambda}$ is constant for these three panels as indicated in the legend of Panel d). For all panels, the cyan line is the M-Köhler curve associated with Eq. (6) and $\beta(N)=3.6\times 10^{-2}\mathrm{s}^{-3/2}$.
The hysteresis behavior pointed out at the end of Multistable Köhler curves is now analyzed when the supersaturation parameter $\lambda$ is allowed to drift through the critical values $\lambda_{c}$ and $\lambda_{h}$, slower compared to the stochastic fluctuations. Such relatively slower variations in $\lambda$ are associated with buoyancy-driven supersaturation changes. Supersaturation $\lambda$ increases as the buoyant parcel rises until the activation threshold $\lambda_{h}$ is attained and some of the contained haze is activated. When the parcel ceases to be buoyant or reaches a temperature inversion, supersaturation decreases to $\lambda_{c}$, below which the whole family of droplets regenerates into haze. During this buoyancy-cycle, the population of particles faces a continuous change in the potential $V_{\lambda}$, as illustrated in Fig. 4, for increasing and decreasing supersaturation levels. It is there shown the asymmetry of activation. Indeed, when a haze particle undergoes activation— starting at point A and following the red, wobbly line—, its deactivation will occur at much lower supersaturation levels— as seen when following the blue wobbly line that ends in point B—. Similar cycles have also been reported in the absence of aerosol curvature and chemistry effects, and uniquely due to noisy supersaturation fluctuations around a subsaturated mean; see (Siewert, Bec, and Krstulovic, 2017, Fig. 4).
Within the present framework, it remains to investigate the effects of (i) different rates of changes in supersaturation and (ii) variations in the magnitude of the fast supersaturation fluctuations as encapsulated by the noise term. The hysteresis phenomenon under the presence of noise has been investigated in previous work Berglund and Gentz (2002), where it is found that a large (additive) noise variance $\sigma$, on average, reduces the area of hysteresis loops. In our context, this implies that droplet activation can occur at much lower supersaturation values compared to $\lambda_{h}$. To numerically examine these ideas, we solve the Eq. (8) associated with the sink term $g(X)=-\beta(N)X^{3/2}$ used in Eq. (6), for specific noise variances and mean supersaturation $\lambda$ ranging from $0.06\%$ to $0.11\%$ with different linear rate $\dot{\lambda}$.
The numerical experiments regarding hysteresis are shown in Fig. 5, where Eq. (6) was integrated using an Euler-Maruyama scheme with a time step of $10^{-2}\mathrm{s}$. In Fig. 5a), b) and c), are shown hysteresis paths solving Eq. (6) for three rates of change $\dot{\lambda}$. As expected, when the rate of change in supersaturation is not slow enough, it can prevent droplet activation— see Fig. 5a)—, whereas sufficiently slow, adiabatic changes enforce the droplet size to follow the M-Köhler curve such as shown in Fig. 5c). As a haze particle experiences an increase in supersaturation— resulting from an adiabatic cooling of a cloud parcel— this will grow by condensation following the Köhler curve, almost regardless of how intense the cooling rate is, here indicated by $\dot{\lambda}$. This is an indication of the stability of the thermodynamic equilibrium of haze particles as opposed to cloud droplets, which are more sensitive the cooling rate $\dot{\lambda}$. Indeed, by comparing Figs. 5b) and c), we observe that only when the cooling rate is slowest do we observe a droplet growth attached to the Köhler curve.
Panels d), e) and f) of Figure 5 provide an illustration of results of Berglund and Gentz (2002) for Eq. (6). Here we show the effects of increasing turbulent noise intensity as a haze particle grows and activates in a slowly cooling air volume. In such case we are in the situation of Fig. 5c) where the particle’s size is dictated by the Köhler curve— modulo zero-mean fluctuations—. As noise variance is increased, the hysteresis loop is collapsed to the extent that droplet activation is expected as soon as $\lambda$ surpasses $\lambda_{c}$. Consequently, it is expected that slowly lifted humid parcels undergoing heaving turbulent updrafts— here encoded in the noise variance— can contain particles that activate at much lower supersaturations compared to that predicted by Köhler.
PI-Chamber Empirical Distributions vs Gibbs States
Small, warm convective clouds are abundant in the atmosphere but hard to measure Koren et al. (2008). Such clouds are often characterized by weak updrafts and small supersaturation, allowing for a significant part of the aerosol to coexist as haze together with the activated part Altaratz et al. (2021); Hirsch et al. (2017). Cloud chambers are often used to study clouds in such thermodynamic conditions experimentally Prabhakaran et al. (2020); Chandrakar et al. (2016).
In such work, the so-called Pi-chamber is seeded with monodisperse families of particles, and turbulent fluctuations are induced by changes in temperature, pressure, and aerosol injection rates. Disregarding transients, the system reaches a steady state in number concentration and size distributions are calculated spanning hazy and active particles. Depending on the supersaturation mean and variance, DSDs display different characteristics involving their multimodality and spread; see (Prabhakaran et al., 2020, Fig. 2). We show here that such experimental distributions are well approximated by the Gibbs states of our Brownian modelling framework once the sink and noise terms are appropriately designed. Our goal is to provide a qualitative and quantitative correspondence between the present theory and the experiments from Prabhakaran et al. (2020).
We first address the sink term design. Note that for a monodisperse family of particles embedded in a cloud parcel, the supersaturation budget must be included in the collective droplet growth Squires (1956); Korolev and Mazin (2003). Because the Pi-chamber is closed and in steady-state, the sink term affecting supersaturation is determined by water vapor condensation onto the particles. The mean sources of supersaturation are taken to be zero, apart from fluctuations encoded in $\sigma(X)$; see Eq. (8). Consequently, and to a first approximation, the initially prescribed supersaturation $\lambda$ is consumed proportionally to the radius of the droplets, so that it evolves according to:
$\lambda(t)\approx\lambda-\beta X(t)^{1/2}+\sigma(X)\dot{W}_{t},$ (13)
where the absorption parameter $\beta$ depends on the number of active droplets, and the effective diffusivity of water vapour in air Khain and Pinsky (2018). By substituting $\lambda$ in the (standard) Köhler equation (1), the new term $-\beta X(t)^{1/2}$ plays the role of a sink term in Eq. (3), so that $g(X)=-\beta X^{1/2}$.
Regarding the term $\sigma(X)\dot{W}_{t}$, the inhomogeneous effect of fluctuations in haze and cloud droplets becomes apparent by qualitatively examining the DSDs obtained in the Pi-chamber; see (Prabhakaran et al., 2020, Fig. 2). To account for such dependence on particle size, we propose the following stochastic parameterization:
$\sigma(X)=\Big{(}\sigma_{1}+\frac{\sigma_{2}-\sigma_{1}}{2}\big{(}1+\psi(X)\big{)}\Big{)},$ (14)
$\psi(X)=\tanh(2\kappa D(X-X^{\ast})).$
Here, the hyperbolic tangent is employed as a step function that allows for interpolating smoothly between the noise strengths that affect haze and activated particles, $\sigma_{1}$ and $\sigma_{2}$ respectively. Thus, we require that $\sigma_{2}>\sigma_{1}>0$ for the reason stated above. The value of the slope is chosen to be sufficiently large— i.e., $102D\kappa=$ 10— so that the hyperbolic tangent approximates a step function, while still exhibiting the attributes of a smooth function, more amenable for analysis and numerical simulations. The “ignition” parameter $X^{\ast}$ is chosen to be equal to $6.2\cdot 10^{-3}\mathrm{s}$ or, in diameter, $d^{\ast}=1.41\mu\mathrm{m}$, which yields a correct location of the haze modes, and it is in the order of the critical Köhler diameter which is $d_{K}=1.7\mu\mathrm{m}$. Such critical value is derived from Eq. (1), as explained in Köhler theory, for the curvature and solubility coefficients $A=1.4\cdot 10^{-3}\mu\mathrm{m}$ and $B=3.5\cdot 10^{-4}\mu\mathrm{m}^{3}$; see Prabhakaran et al. (2020).
We, thus, arrive at the following Brownian model to describe cloud droplet size evolution in the Pi-chamber:
$\dot{X}=\lambda-f(X)-\beta X^{1/2}+\sigma(X)\dot{W}_{t},$ (15)
with $\sigma$ given by (LABEL:eta_term_chamber), and we recall that $f$ is given by (2).
By solving Eq. (11) associated with Eq. (15), we can derive a relationship (valid for $X$ sufficiently bigger than $X^{\ast}$) determining the parameter $\beta$ in terms of the modal size $X_{c}$ of the activated mode:
$\beta=\frac{1}{X_{c}^{2}}\left(\lambda X_{c}^{3/2}-\frac{A}{(2D)^{1/2}}X_{c}+\frac{B}{(2D)^{3/2}}\right).$ (16)
Eq. (16) plays a key role to estimate $\beta$: once $\beta$ is known, the only parameters left for calibration of the model are $\sigma_{1}$, $\sigma_{2}$ and the stiffness parameter $\kappa$.
Figure 6: Panels a), b) and c): Köhler and M-Köhler curves (dashed and solid curves, respectively) for each supersaturation value considered in PI-Chamber Empirical Distribution vs Gibbs States (Cases I, II, and III). The parameters of the Köhler curve are $A=1.4\cdot 10^{-3}\mathrm{\mu m}$ and $B=3.5\cdot 10^{-4}\mathrm{\mu m^{3}}$ and the other parameters of model Eq. (15) are given in Table 2. The black dot denotes the location of the metastable state induced by the state-dependent noise; see Noise-induced metastability in Methods. Panels d), e) and f): Experimental DSDs from Prabhakaran et al. (2020) vs Gibbs states from Eq. (9). Both the data histograms and Gibbs states are normalized according to their integral. The insets show the transformed potentials (due to Lamperti transformation) explaining the shapes of the Gibbs states; see The Lamperti transformation in Methods.
We now confront the ability of our Brownian model (15) to approximate experimental DSDs by the Gibbs states formula (9). Three experimental setups are considered in that respect, following those of the Pi-chamber experiments from Prabhakaran et al. (2020) corresponding to different supersaturation mean and fluctuation properties reported in such paper. These are referred to as according to Prabhakaran et al. (2020): a) mean-dominated, b) fluctuation influenced and c) fluctuation dominated acivation.
Within our modeling framework, these regimes can be organized as follows:
  1. Case I: $\lambda>\lambda_{K}\rightarrow$ mean-dominated activation
  2. Case II: $\lambda_{c}<\lambda<\lambda_{K}\rightarrow$ fluctuation influenced activation
  3. Case III: $\lambda<1\rightarrow$ fluctuation dominated activation
The parameter $\lambda_{K}$ is the critical Köhler threshold, which is derived to be $\lambda\approx 1.00108$, for the type of aerosols considered in Prabhakaran et al. (2020). The steady-state supersaturation value in each case is also derived from Prabhakaran et al. (2020). Their theoretical calculations reveal indeed that Case I corresponds to $\lambda=1\%$. While the value for Case II is not explicitly provided, it can be inferred from our framework that $\lambda$ is close to $\lambda_{h}$ from below since a bimodal DSD is observed, with high peaks in the haze and activated domain: $\lambda=0.1\%$ is taken for this case. Finally, Case III is a subsaturated regime ($\lambda<0$) and as such we choose $\lambda$ to be given by the Köhler’s value $f(X)$ at the haze peak’s location $X=X_{h}$, this gives $\lambda=-1\%$.
With these parameter values for $\lambda$, the parameter $\beta$ can be estimated for Cases I and II according to Eq. (16), in which $X_{c}$ is chosen to be the experimental DSD peak’s location that corresponds to activated droplets. For Case I, the modal activated diameter is $d_{c}=18.109\mu\mathrm{m}$, giving $X_{c}=1.025\mathrm{s}$ and $\beta=9.7\cdot 10^{-3}\mathrm{s}^{-1/2}$. In Case II, the modal activated diameter is $d_{c}=9.141\mu\mathrm{m}$, giving $X_{c}=0.2611\mathrm{s}$ and $\beta=1.4\cdot 10^{-3}\mathrm{s}^{-1/2}$. These modal values are obtained from the data of Prabhakaran et al. (2020). For Case III, though, since the system is subsaturated and haze particles are more numerous—see Table 1 of Prabhakaran et al. (2020)—the term $\beta$ is practically zero. The corresponding original Köhler and M-Köhler curves are shown in Fig. 6a) and b), as dashed and solid curves, respectively, and shown as functions of the diameter $d=2\sqrt{2DX}$. For Case III, because there is not sink term, only the original Köhler curve is shown.
Once the parameters for the relative disturbances, $\sigma_{1}$, $\sigma_{2}$ and $\kappa$ are chosen as listed in Table 2, the corresponding Gibbs states $\rho_{\lambda}$ given by Eq. (9) exhibit striking skills in reproducing the empirical DSDs in each of the three Pi-chamber experiments taken from Prabhakaran et al. (2020). The results are shown in the Figs. 6d), e) and f). The shapes of the Gibbs states are highly correlated to those of the transformed potentials shown in the insets of Fig. 6, obtained by the Lamperti transformation; see Methods. We observe in particular that the potential wells’ locations match the Gibbs states modes’ locations. Recall that the latter transformation allows for identifying the effective potential for the stochastic model (15) subject to size-dependent stochastic disturbances such as given by (LABEL:eta_term_chamber). With this effective potential, Eq. (15) can be interpreted as a Brownian particle evolving within that potential whose fluctuations are just driven by a white noise, independently of the size of the particle.
As sanity check, one observes that the experimental DSDs’ modes are well predicted by Eq. (11), as shown in Fig. 7 of Noise-induced metastability, especially in Cases I and II, where the classical Köhler theory is unable to predict an activated mode, since $\lambda>\lambda_{K}$. Noteworthy is the creation of a metastable haze-state in Case I resulting from the interaction between the noise term $\sigma(X)\dot{W}_{t}$ and the nonlinear drift terms $-f(X)-\beta X^{1/2}$ as revealed by the transformed potential shown in the inset of Figs. 6d); see Noise-induced metastability in Methods.
Table 2: Supersaturation $\lambda$ and parameter values of the sink term and state-dependent noise $\sigma(X)$ in Eq. (LABEL:eta_term_chamber).
[…]
Case I $1$ $9.6\cdot 10^{-3}$ $3.75\cdot 10^{-2}$ $6.25\cdot 10^{-2}$ 1.41 10
Case II $0.1$ $1.4\cdot 10^{-3}$ $7.5\cdot 10^{-3}$ $1.5\cdot 10^{-2}$ 1.41 10
Case III $-1$ 00 $5\cdot 10^{-3}$ $1.5\cdot 10^{-2}$ 1.41 10
DISCUSSION
Köhler theory stands in the foundation of cloud physics. For a given dry aerosol size and composition, it predicts the particle’s state, haze, or activated droplet as a function of the particle size and the supersaturation. The Köhler curve (Fig. 1a), has two fixed-point regimes bounded by the curve’s mode. The curve left to the mode describes a stable fixed point, whereas any particle that resides on the curve right to the mode is on an unstable fixed point such that a perturbation to the right (or upward) will shift the particle to a continuous growth by diffusion and a perturbation to the left (or downward) will send the particle to the haze state. Warm and small clouds that are driven by weak temperature or humidity perturbations Eytan et al. (2020); Altaratz et al. (2021); Hirsch et al. (2017) are superabundant and have an important radiative effect. However, due to their sizes, weak optical signature, and short lifetimes, these clouds are mostly overlooked, and their properties are poorly understood. Such clouds often exist near the transition between haze to cloud Koren et al. (2009) a regime that can be studied in cloud chambers Prabhakaran et al. (2020); Chandrakar et al. (2016).
Here, we address these clouds by extending Köhler theory to population of droplets embedded in turbulent environments. While activation is explained by Köhler’s theory, the reverse process, i.e. deactivation, is dynamically different. It has been documented in previous works Arabas and Shima (2017); Korolev and Mazin (2003) and in the present study that deactivation is intimately tied to the hysteresis phenomenon through which deactivation occurs at a smaller supersaturation compared to the critical Köhler value $\lambda_{K}$. The addition of a sink term in the condensational growth equation— see Eq. (3)— yields hysteresis directly from an analytical perspective: in order to deactivate a family of monodisperse cloud droplets, the supersaturation needs to be brought strictly below the critical value $\lambda_{K}$ predicted by Köhler, see Figs. 1b). The origin of this behavior is the presence of another saddle-node bifurcation that results from the bending of the M-Köhler curve introduced in Multistable Köhler curves This bending translates further into the appearance of another stable equilibrium on the activated droplet spectrum as a balance between condensational instability and sink factors, like supersaturation consumption or particle removal. We refer to the section Multistability in Methods for conditions on the sink function $g$ ensuring for such a bending to occur, and thus the creation of an extra saddle-node bifurcation in Eq. (3).
The introduction of Brownian noise in the condensational growth equation parameterizes the influence of supersaturation fluctuations and gives analytical formulas for computing DSDs, here identified as Gibbs states; see Eq. (9). Such probability distribution is a general formulation of DSDs found in the literature McGraw and Liu (2006); Chandrakar et al. (2020); Krueger (2020), with the addition of size dependent fluctuations. The resulting DSDs are classified into three types according to the mean supersaturation parameter $\lambda$. The first corresponds to the case where supersaturation fluctuations do not have the strength to activate droplets for a prolonged period of time and, hence, the size distribution is unimodal around the haze state. In the second case, activation is allowed and if the threshold $\lambda_{h}$ is surpassed, haze particles cease to be stable and a single DSD mode is observed peaking over the activated sizes. Consequently, even if Köhler’s threshold supersaturation is not attained, there is a finite probability that particles activate and remain so stably. The associated timescales for activation is therefore related to the average time it takes to droplets to gain sufficient energy to surpass the potential barrier of activation— see Fig. 1c)—, and they are given precisely by Kramers’ formula Eq. (12). In fact, when a cloud parcel is adiabatically cooled, supersaturation fluctuates around a slowly increasing mean. Neglecting noisy fluctuations would incur into an overestimation of the superstaruation threshold for activation, as shown in every panel of Fig. 5, where the interplay between noise and activation hysteresis is illustrated. It is realized that on an activation-deactivation cycle, on average, activation occurs at a much lower supersaturations compared to $\lambda_{h}$, and the larger the variance the lower such threshold is.
The confrontation of our population Köhler scheme within our Brownian modeling framework to steady-state cloud experiments conducted in the Pi-chamber has revealed a striking consistency. Particularly at capturing to a high degree of fidelity the complexity of the DSDs’ shapes along with their bimodal features as reported in the experimental study of Prabhakaran et al. (2020), for different saturation regimes. Under such a setting, it remains to investigate the slow adiabatic changes in mean supersaturation. Based on the analytical insights gained on the droplet activation-deactivation hysteresis (see Fig. 5), we foresee that the same hysteresis behavior should also be observed in the experimental framework of Prabhakaran et al. (2020).
Weak clouds in nature are likely to be exposed to aerosols of several sizes and compositions. However, the weak supersaturation may act as a selective filter, making the largest haze bins the leading players, yielding conditions that are not so different from our theoretical setting. Moreover, the additional Brownian noise blurs the particles’ homogeneity due to the monodisperse nature of the initial conditions. Nevertheless, cases with polydisperse particles of different chemical natures and, hence, different Köhler curves are planned to be investigated.
METHODS
Multistability. In order for model (3) to allow for multiple stable equilibria, it is enough if the the M-Köhler curve $f-g$ has at least a local minimum and a local maximum. This property is to guarantee that the zeros of $\lambda-(f-g)$ correspond to at least two stable equilibria and an unstable one, for a range of values of $\lambda$.
To simplify the presentation of the results stated in Proposition 1 below, we introduce the following parameters:
$\tilde{A}$ $=\frac{A}{(2D)^{1/2}},$ (17a)
$\tilde{B}$ $=\frac{B}{(2D)^{3/2}}.$ (17b)
We then have the following sufficiency result for the existence of a local maximum and a local minimum:
Proposition 1.
Let $f$ be the function given in Eq. (2) and let $g:\mathbb{R}^{+}\longrightarrow\mathbb{R}^{-}$ be a function of the form $g(X)=-\beta X^{\alpha}$, where $\beta>0$ and $\alpha>1$. If $\dot{g}(5\tilde{B}/\tilde{A})>\dot{f}(5\tilde{B}/\tilde{A})$, the M-Köhler curve in Eq. (3) has at least a local maximum and a local minimum.
Proof.
We need to show that the function $\dot{f}-\dot{g}$ has two zeros, and thus it suffices to prove that $\dot{f}-\dot{g}$ has at least two changes of sign. This function reads, for $X>0$, as:
$\dot{f}(X)-\dot{g}(X)=-\frac{\tilde{A}}{2}X^{-3/2}+\frac{3\tilde{B}}{2}X^{-5/2}+\alpha\beta X^{\alpha-1}.$ (18)
When $\alpha=3/2$ as in Eq. (6), one recovers that finding the zeros of $\dot{f}-\dot{g}$ is equivalent to finding the roots of the polynomial equation (7).
First of all, $\dot{f}$ has a single local and global minimum. Indeed, we see this by examining $\ddot{f}$:
$\ddot{f}(X)=\frac{3}{4}\tilde{A}X^{-5/2}-\frac{15}{4}\tilde{B}X^{-7/2},$ (19)
which has a unique zero at $X=5\tilde{B}/\tilde{A}$, and
$\dot{f}\left(\frac{5\tilde{B}}{\tilde{A}}\right)=-\frac{\tilde{A}}{5}\left(\frac{\tilde{A}}{5\tilde{B}}\right)^{\frac{3}{2}}<0.$ (20)
Furthermore, the two limits below hold:
$\lim_{X\rightarrow 0^{+}}\dot{f}(X)=+\infty$ (21a)
$\lim_{X\rightarrow\infty}\dot{f}(X)=0.$ (21b)
These two limits together with Eq. (20) indicate that $X=5\tilde{B}/\tilde{A}$ is a global minimum.
Because of the smoothness of $f$ and $g$ on $\mathbb{R}^{+}$, the limit in Eq. (21a), and— by hypothesis— $\alpha>1$, we have:
$\lim_{X\rightarrow 0^{+}}\dot{f}(X)-\dot{g}(X)=+\infty,$ (22)
there exists $X_{1}>0$ such that $\dot{f}(X_{1})-\dot{g}(X_{1})>0$. Moreover, by hypothesis, $\dot{f}(5\tilde{B}/\tilde{A})-\dot{g}(5\tilde{B}/\tilde{A})<0$, and that gives the first change of sign of the function $\dot{f}-\dot{g}$.
Lastly, by assumption we have that $\alpha>1$, so the following limit holds
$\lim_{X\rightarrow\infty}\dot{g}(X)=\lim_{X\rightarrow\infty}-\alpha\beta X^{\alpha-1}=-\infty.$ (23)
Considering this limit and the one in Eq. (21b), there exists $X_{2}>5\tilde{B}/\tilde{A}$ such that $\dot{f}(X_{2})-\dot{g}(X_{2})>0$. This gives the second change of sign.
Due to the continuity and smoothness of the functions $f$ and $g$ on $\mathrm{R}^{+}$, the function $\dot{f}-\dot{g}$ must have at least two zeros. ∎
The confining potential. In order for a gradient system to possess a Gibbs state, it is necessary that the associated potential is confining see e.g. (Pavliotis, 2014, §4). We provide here sufficient mathematical conditions by which a sink function $g$ in Eq. (8) ensures the confining property. Generally, it amounts to imposing a sufficiently strong decay rate to the sink term.
The potential function $V_{\lambda}$ is said to be confining if:
$\lim_{X\rightarrow 0^{+}}V_{\lambda}(X)=\lim_{X\rightarrow\infty}V_{\lambda}(X)=\infty,$ (24)
and if for any $\xi>0$,
$\int_{0}^{\infty}e^{-\xi V_{\lambda}(X)}\,\mathrm{d}X<\infty.$ (25)
Below, we provide a class of sink functions for which the associated M-Köhler curve has a confining potential.
Proposition 2.
Let $\lambda$ be in $\mathbb{R}$, $f$ be as in Eq. (2) and let $g:\mathbb{R}^{+}\longrightarrow\mathbb{R}^{-}$ be defined as $g(X)=-\beta X^{\alpha}$, for $X,\alpha,\beta>0$. Then, the potential $V_{\lambda}$ associated with Eq. (8) is confining.
Proof.
By the definition of $f$, $g$ and the potential $V_{\lambda}$, we find that:
$-V$ (${}_{\lambda}(X)=\int^{X}\left(\lambda-f(X)+g(X)\right)\,\mathrm{d}X$ (26a)
$=\lambda X-2\tilde{A}X^{1/2}-2\tilde{B}X^{-1/2}-\frac{\beta}{1+\alpha}X^{1+\alpha},$ (26b)
where the parameters $\tilde{A}$ and $\tilde{B}$ are defined in Eq. (17). Then,
$\lim_{X\rightarrow 0^{+}}V_{\lambda}(X)=\lim_{X\rightarrow 0^{+}}2\tilde{B}X^{-1/2}=\infty,$ (27a)
$\lim_{X\rightarrow\infty}V_{\lambda}(X)=\lim_{X\rightarrow\infty}\frac{\beta}{1+\alpha}X^{1+\alpha}=\infty.$ (27b)
We are left with showing the integrability condition (25). The integral of the exponential of the potential satisfies, for any $\xi>0$, the following inequalities:
$\int_{0}^{\infty}e^{-\xi V_{\lambda}(X)}\,\mathrm{d}X$ (28a)
$=\int_{0}^{\infty}e^{\xi\left(\lambda X-2\tilde{A}X^{1/2}-2\tilde{B}X^{-1/2}-\frac{\beta}{1+\alpha}X^{1+\alpha}\right)}\,\mathrm{d}X$ (28b)
$\leq\int_{0}^{\infty}e^{\xi\left(\lambda X-\frac{\beta}{1+\alpha}X^{1+\alpha}\right)}\,\mathrm{d}X$ (28c)
$=\int_{0}^{\infty}e^{-X\left(\frac{\xi\beta}{1+\alpha}X^{\alpha}-\xi\lambda\right)}\,\mathrm{d}X$ (28d)
$\leq\int_{0}^{X_{0}}e^{-X\left(\frac{\xi\beta}{1+\alpha}X^{\alpha}-\xi\lambda\right)}\,\mathrm{d}X+\int_{X_{0}}^{\infty}e^{-X}\,\mathrm{d}X<\infty,$ (28e)
for any $\xi>0$ and where $X_{0}=((1+\xi\lambda)(1+\alpha)/\xi\beta)^{1/\alpha}$ is such that:
$\frac{\xi\beta}{1+\alpha}X_{0}^{\alpha}-\xi\lambda=1.$ (29)
This proves the confining property. ∎
Hysteresis path algorithm. To produce the hysteresis paths shown in Fig. 5, we solve Eq. (6) for slowly varying values of the supersaturation parameter $\lambda$ and Brownian noise’s parameter $\sigma$ taken to be additive, i.e. $\sigma(X)=\sqrt{2\varepsilon}$. The slow drift of the supersaturation parameter is organized as follows. We start by dividing an interval $[\lambda_{0},\lambda_{N+1}]\supset(\lambda_{c},\lambda_{h})$ uniformly into $N+2$ grid points $\{\lambda_{j}\}_{j=0}^{N+1}$ such that $\lambda_{j}=\lambda_{0}+j\delta\lambda$, with $\delta\lambda=(\lambda_{h}-\lambda_{c})/(N-1)$ and $\lambda_{0}=\lambda_{c}-\delta\lambda/2$. Note that $\lambda_{N+1}>\lambda_{h}$ and $\lambda_{N}<\lambda_{h}$.
To form the lower branch of this hysteresis path, we start with $\lambda=\lambda_{0}$ and from the haze state $X_{h}(\lambda_{0})$, and solve Eq. (6) using an Euler-Maruyama scheme (Kloeden and Platen, , Chapter 9.1) for one timestep; $\mathrm{d}t=1$. Then $\lambda$ is updated to $\lambda_{0}+j\delta\lambda$ and Eq. (6) is iterated over another timestep by still using Euler-Maruyama. The process is repeated $N+1$ times until $\lambda_{N+1}>\lambda_{h}$. The reverse trajectory is obtained by repeating the algorithm, although starting with $\lambda=\lambda_{N+1}$ at $X_{c}(\lambda_{N+1})$. More precisions are given in Algorithm 1 and hereafter.
Algorithm 1 Hysteresis path computation
1:  $\delta t$ and $N$ is prescribed by the user
2:  $\delta\lambda=(\lambda_{h}-\lambda_{c})/(N-1)$
3:  $\lambda^{+}_{vals}=\{\lambda_{0}+j\delta\lambda\}_{j=0}^{N+1}$, with $\lambda_{0}=\lambda_{c}-\delta\lambda/2$
4:  $\lambda^{-}_{vals}=\{\lambda_{N+1}-j\delta\lambda\}_{j=0}^{N+1}$
5:  Form the set $\lambda_{vals}=\{\lambda^{+}_{vals},\lambda^{-}_{vals}\}$
6:  Let $X$ be a $2(N+2)$-dimensional vector approximating the hysteresis path
7:  We choose the initial datum to be $X_{h}\left(\lambda_{0}\right)$, i.e. $X(1)=X_{h}\left(\lambda_{0}\right)$
8:  for $j=0:2N+3$ do
9:     $\mu=\lambda_{vals}(j+1)$
[…]
11:     $X(j+1)=X(j)-\delta tV_{\mu}(X(j))+\sqrt{2\varepsilon\delta t}\zeta_{j}$, where $\zeta_{\ell}$ are independent normally distributed random variables with mean zero and variance unity
12:  end for
The key parameter to produce panels a), b) and c) of Fig. 5 is $\dot{\lambda}$. Note that according to the previous algorithm, $\dot{\lambda}\approx\delta\lambda/\delta t$. Consequently, the larger $N$ is, the smaller $\delta\lambda$, and vice versa.
Figure 7: Noise-induced metastable states. Panels a), b) and c), correspond to the three different supersaturation configurations as indicated in the titles and corresponding to those of Fig.  6. The blue curves show the growth rate function $-V^{\prime}_{\lambda}$ in Eq. (8), such as associated with the M-Köhler curve $f(X)+\beta X^{1/2}$. The red curves show the derivative of the noise covariance after division by a factor four corresponding to the RHS of Eq. (34). The vertical dashed lines indicate the location of the solutions to Eq. (34) which also correspond to the haze, unstable and activated metastable states, denoted by $d_{h},d_{u}$ and $d_{c}$, respectively.
The Lamperti transformation. For scalar stochastic differential model, it is possible to map a stochastic differential equation (SDE) driven by state-dependent noise onto an SDE driven by an additive noise. We recall this basic idea. Consider the following scalar SDE
$\dot{X}=-V(X)+\sigma(X)\dot{W}_{t},$ (30)
with $V$ and $\sigma$ smooth functions, such that $\sigma>0$. Consider the change of variables, $h(X)=\int^{X}\frac{1}{\sigma(X)}\,\mathrm{d}X$, then by application of Itô’s formula, the original equation (30) is transformed into
$\dot{Y}=-\widehat{V}(Y)+\dot{W}_{t},$ (31)
with
$\widehat{V}(Y)=\frac{V(h^{-1}(Y))}{\sigma(h^{-1}(Y))}-\frac{1}{2}\sigma^{\prime}(h^{-1}(Y)).$ (32)
This is called the Lamperti transformation (Pavliotis, 2014, §3.6).
The Lamperti transformation allows for identifying the effective potential, $\widehat{V}$, for the stochastic model (30) subject to state-dependent stochastic disturbances given by $\sigma(X)\dot{W}_{t}$. With this effective potential, Eq. (30) can be interpreted as a Brownian particle evolving within that potential whose fluctuations are just driven by a white noise, independently of the state $X$. In particular even if the original potential $V$ does not exhibit a well over a certain range of $X$-values, the transformed potential can exhibit such local minimum. We talk then of noise-induced metastability in case of creation of an extra metastable state resulting from the interaction of the drift term with the diffusion term $\sigma(X)\dot{W}_{t}$.
Noise-induced metastability. By applying (32) to our stochastic droplet model (15) in Pi-chamber, we obtain:
$\widehat{V}^{\prime}_{\lambda}(Y)=-\frac{V^{\prime}_{\lambda}(X)}{\sigma(X)}-\frac{1}{2}\sigma^{\prime}(X),\mbox{ with }Y=\int^{X}\sigma^{-1}(X)\,\mathrm{d}X,$ (33)
in which $V_{\lambda}^{\prime}(X)=-\lambda+f(X)+\beta X^{-1/2}.$
Note that the appearance of equilibria non-trivially depends on both the potential $V_{\lambda}$ and the noise function $\sigma$. Indeed, as a consequence of Eq. (33), the metastable states are explicitly given by solving $\widehat{V}^{\prime}_{\lambda}(Y)=0$, namely by solving:
$-V^{\prime}_{\lambda}(X)=\frac{1}{2}\sigma^{\prime}(X)\sigma(X),$ (34)
recovering this way Eq. (11). Thus, if one wishes to match the locations of these metastable states with let us say the modes of an empirical distribution (as in PI-Chamber Empirical Distributions vs Gibbs States), one needs to solve Eq. (34). As a consequence, this equation imposes constraints on the choice of the parameters involved in $\sigma$ and $V_{\lambda}$.
More precisely, we have that (34) rewrites as
$\lambda-\tilde{A}X^{-1/2}+\tilde{B}X^{-3/2}-\beta X^{1/2}=\frac{1}{2}\sigma^{\prime}(X)\sigma(X),$ (35)
where $\tilde{A}$ and $\tilde{B}$ are defined in Eq. (17). By observing that $\sigma(X)=\sigma_{1}+\frac{\sigma_{2}-\sigma_{1}}{2}\big{(}1+\tanh(\kappa(X-X^{\ast}))\big{)}\longrightarrow 0$ as $X\rightarrow\infty$, one can get then, when $\kappa\gg 1$ and $X$ is sufficiently larger than $X^{\ast}$, good approximations of the solutions $X$ to Eq. (35) by solving:
$\lambda X^{3/2}-\tilde{A}X+\tilde{B}-\beta X^{2}=0,$ (36)
We adopt this strategy for Cases I and II of PI-Chamber Empirical Distribution vs Gibbs States.
The sink parameter $\beta$ has a direct influence on the location of the metastable states and vice versa. One can invert equations (36) to find the value of $\beta$ in terms of the activated mode $X_{c}$:
$\beta=\frac{1}{X_{c}^{2}}\left(\lambda X_{c}^{3/2}-\tilde{A}X_{c}+\tilde{B}\right),$ (37)
leading thus to Eq. (16).
The blue and red curves in Fig. 7 correspond to the left-hand-side (LHS) and right-hand-side (RHS) of Eq. (34), respectively, for the three case studies of PI-Chamber Empirical Distributions vs Gibbs States. The intersection of these two curves determines the location of the metastable states indicated by the dashed vertical lines for Cases I, II and III with the parameter configuration shown in Table 2.
Acknowledgements This work has been partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 810370). MSG is grateful to the Feinberg Graduate School for their support through the Dean of Faculty Fellowship.
Author contributions: MDC conceived the presented idea. MSG led the analyses and MDC supported. MSG, MDC, OA, and IK discussed the results and wrote the manuscript. All co-authors provided critical feedback and helped shape the research, analysis and manuscript.
Competing interests: The authors declare no competing interests.
Data availability: All experimental data used in this study are based on the article of Prabhakaran et al. (2020) and are publicly available at: https://digitalcommons.mtu.edu/data-files/3/. The numerical data obtained for this article are available upon request to the authors.

### v2

@@
 Gibbs states and Brownian models for coexisting haze and cloud droplets
-Abstract
-Clouds microphysics describes the formation and evolution of cloud droplets, rain, and ice particles.
-It is among the most critical factors in determining the cloud’s size, lifetime, precipitation, and radiative effect.
-Among all cloud types, the small clouds, characterized by weak updrafts, that are close to the haze-to-cloud transition pose challenges in measuring them and understanding their properties.
-They are superabundant but hard to capture by satellites and often falsely regarded as aerosols.
-Köhler’s theory explains droplet activation and their growth by condensation at the earliest stages of cloud development.
-It fully describes the thermodynamic state of a single drop but falls short when explaining the collective behavior of large populations of particles.
-This is especially important when the supersaturation pool is limited.
-We present an analytical framework to extend Köhler’s theory to coexisting cloud droplets.
-Our results suggest hysteresis and asymmetry in the process of droplet activation and deactivation.
-The turbulent nature of clouds is incorporated into our model formulation as Brownian noise to provide explicit droplet size distributions and activation timescales.
-The theoretical findings are confronted with experimental data stemming from laboratory clouds created in the Pi convection chamber, suggesting a new way of understanding haze-to-cloud transitions and small cloud formation processes.
+Abstract.
+Cloud microphysics studies include how tiny cloud droplets grow, and become rain.
+This is crucial for understanding cloud properties like size, lifespan, and impact on climate through radiative effects.
+Small, weak-updraft clouds near the haze-to-cloud transition are especially difficult to measure and understand.
+They are abundant but hard to capture by satellites.
+Köhler’s theory explains initial droplet growth but struggles with large particle groups.
+Here, we present a stochastic, analytical framework building on Köhler’s theory to account for (monodisperse) aerosols and cloud droplets interaction through competitive growth in a limited water vapor field.
+These interactions are modeled by sink terms while fluctuations in supersaturation affecting droplet growth are modeled by nonlinear, white noise terms.
+Our results identify hysteresis mechanisms in the droplet activation and deactivation processes.
+Our approach allows for multimodal cloud’s droplet size distributions supported by lab experiments, offering a new perspective on haze-to-cloud transition and small cloud formation.
+Keywords. Köhler Theory |||| Droplet Size Distributions |||| Stochastic Modeling |||| Hysteresis |||| Gibbs States
 Introduction
-Clouds play a key role in the climate energy balances and in the water cycle Bengtsson (2010).
-A large part of the uncertainties in climate predictions is attributed to limitations in our understanding of cloud physics and therefore, how to parameterize such clouds in global climate models Zelinka et al.
-(2020).
-Out of all cloud types, a great part of the cloud research challenges is attributed to small, warm clouds that are highly abounded but hard to measure.
-Many of these clouds are too small relative to earth-observing satellites’ resolution, and/or their optical signature is too weak Koren et al.
-(2007); Várnai and Marshak (2009); Eytan et al.
-(2020).
+Clouds play a key role in the climate energy balances and in the water cycle [Ben10].
+A large part of the uncertainties in climate predictions is attributed to limitations in our understanding of cloud physics and therefore, how to parameterize such clouds in global climate models [ZMM+20].
+Out of all cloud types, small, warm clouds, are highly abundant and pose serious challenges for cloud research.
+Their diminutive size often falls below the resolution limits of Earth-observing satellites, making direct observation difficult.
+Additionally, their weak optical signatures can hinder detection and measurement [KRK+07, VM09, EKA+20].
+These factors contribute to a substantial knowledge gap regarding the role and impact of small, warm clouds in the global climate system.
 The cloud’s droplet size distribution (DSD) is among the most important microphysical properties.
 The DSD properties are coupled to all dynamical, turbulent, optical, and stochastic cloud processes.
-DSDs affect the cloud’s evolution in time, rain processes, cloud depth, size, and lifetime Khain and Pinsky (2018).
+DSDs affect the cloud’s evolution in time, rain processes, cloud depth, size, and lifetime [KP18].
 When supersaturation is high enough, cloud condensation nuclei (CCNs) are activated into cloud droplets and grow by condensation.
 This is the main mechanism for droplet growth at the early stages of a warm cloud.
-If the cloud’s dynamics is weak (shallow clouds) or if the aerosol concentration is high, growth droplets by condensation could be the main mechanism throughout the whole cloud’s lifetime as droplets do not become large and varied enough to support efficient autoconversion of droplets by the stochastic processes of collision and coalescence (Rogers and Yau, 1989, §7-8).
-In warm clouds that are controlled by weak updrafts and low supersaturations, only a small part of the aerosol (the large and most hygroscopic) are activated.
-A competition between the droplet curvature (Kelvin) effect, which suppresses activation, and the solute (Raoult) effect, which reduces the critical supersaturation for activation, Köhler (1936) allows for suspended particles in air parcels to coexist in two main thermodynamical states known as: haze and activated droplet Khain and Pinsky (2018).
-While hazy particles are in stable equilibrium, activated particles harvest the available water vapor and grow until equilibrium is restored.
+If the cloud’s dynamics is weak (shallow clouds) or if the aerosol concentration is high, condensation could be the main growth mechanism throughout the whole cloud’s lifetime as droplets do not become large and varied enough to support efficient autoconversion of droplets by the stochastic processes of collision and coalescence; see Chapters 7 and 8 in [RY89].
+Köhler theory provides a framework for understanding the activation of aerosol particles into cloud droplets.
+In low supersaturation conditions, only a small subset of the aerosol population, typically the largest and most hygroscopic particles, are activated.
+The activation process is governed by a delicate balance between the Kelvin effect, which opposes activation due to the curvature of the droplet surface, and the Raoult effect, which favors activation due to the presence of dissolved solutes.
+This interplay is central to Köhler theory [Köh36].
+As a result of this competition, aerosol particles can exist in two distinct thermodynamic states: haze particles and activated droplets [KP18].
+Haze particles remain typically in stable equilibrium with the surrounding environment, while activated droplets grow by diffusion.
+However other microphysics regimes in a turbulent flow can lead to intricate interactions between haze and activated droplets, resulting e.g.
+into nonlinear behaviors such as oscillatory pulses of activation at low particle concentrations [YSS+24].
 The mechanisms of activation and deactivation are, however, not symmetric.
 In order to activate a cloud droplet, it is necessary to attain a critical level of supersaturation $\lambda_{K}$, typically well above $0\%$.
-When supersaturation is decreased to $\lambda_{K}$ in the deactivation process, an activated cloud droplet is big enough to still experience diffusion of water vapor onto its surface and, hence, it is expected that this positive feedback makes the threshold for deactivation strictly lower than that of $\lambda_{K}$.
-Consequently, activation-deactivation undergoes hysteresis, which has been previously reported in the cloud physics literature Korolev and Mazin (2003); Arabas and Shima (2017).
-From a theoretical perspective, in the work of Arabas and Shima (2017) it was shown on analytical grounds that the activation process of a cloud particle undergoes a saddle-node bifurcation and, when coupled to supersaturation evolution, hysteresis is numerically observed in a parcel model and attributed to kinetic limitations and the presence of a cusp bifurcation (Kuznetsov, 2004, §8.2).
-Another level of complexity in the determination of microphysical properties concerns clouds living in eminently turbulent environments.
-Indeed, turbulence heavily influences the life cycle of clouds; see Shaw (2003) for a review on the topic.
-In particular, the incorporation of turbulent models provides the means to explain droplet size spectra broadening Sardina et al.
-(2015); Grabowski and Abade (2017), as well as describe the mixing of the cumulus clouds with their exterior Abade, Grabowski, and Pawlowska (2018); Eytan et al.
-(2020) and parameterize mixed-phase clouds in global circulation models Field et al.
-(2014); Furtado et al.
-(2016).
-In the context of warm clouds, localized temperature gradients and turbulence at the smallest scales provoke supersaturation fluctuations that allow activation of haze particles even in mean-subsaturated environments Prabhakaran et al.
-(2020), and similarly, cloud particles susceptible of deactivating Abade, Grabowski, and Pawlowska (2018).
-Thus, microphysical fluctuations are a key feature in obtaining faithful descriptions of droplet size spectra.
-This research explores the activation and deactivation of cloud droplets in warm clouds, offering new theoretical perspectives.
-We aim to find analytical formulas for describing the size distribution of droplets (DSDs) when both small cloud droplets and haze particles coexist, assuming condensation as the only growth mechanism.
-Similar to previous work Arabas and Shima (2017), we analyze how droplet growth due to condensation changes dramatically as the surrounding air’s supersaturation level varies.
-To the Köhler instability we add a sink term in growth rate which aims at (i) modeling the supersaturation budget, and (ii) parameterising other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al.
-(2014); Altaratz et al.
-(2021).
-The resulting equation describing condensation growth for droplets of a uniform size allows for multiple stable states to co-exist besides the ”haze” state, meaning the system exhibits multistability.
-Although the haze state remains stable, this multistability suggests that activated droplets can get stuck around a specific size due to the balancing act between two factors: the Köhler instability and mechanisms that decrease supersaturation or remove particles; see also Korolev and Mazin (2003); Arabas and Shima (2017); Chandrakar et al.
-(2020).
-However, turbulence effects throw a wrench in the picture.
+When the supersaturation is decreased to $\lambda_{K}$ during the deactivation process, an activated cloud droplet may remain sufficiently large to continue experiencing diffusion of water vapor onto its surface.
+This positive feedback mechanism results into a deactivation threshold that is strictly lower than $\lambda_{K}$.
+In this case, activation-deactivation exhibits hysteresis, a phenomenon previously documented in the cloud physics literature [KM03, AS17].
+From a nonlinear dynamics perspective, such an hysteresis phenomenon results from the presence of two saddle-node bifurcation curves that meet tangentially (Chapter 8.2 in [Kuz04]).
+Recall that a saddle-node bifurcation is a type of bifurcation where two equilibrium points “collide” and disappear as a system’s parameter is varied (here supersaturation).
+It is at the core of the notion tipping points [AWVC12] and occurs in many fields of physics such as combustion theory [BE13, Che18], energy balance climate models [Ghi76, NCCJ81], or ocean dynamics [HSA+11, WCD+19, CLM23, LC23].
+In the context of cloud physics, the authors in [AS17] have shown that the activation process of a cloud particle following the Köhler condensational growth equation, undergoes such a saddle-node bifurcation in which the haze equilibrium merges with an unstable equilibrium.
+At this bifurcation point, the stability of the haze equilibrium is lost due to this merging.
+On analytical and numerical grounds it was shown in [AS17] that hysteresis does occur for air-parcel models accounting for temperature and pressure variations, suggesting the presence of the other saddle-node bifurcation where the activated equilibrium merges with an unstable one.
+In this work, we provide a general framework for hysteresis to take place from the Köhler condensational growth equation alone by incorporating a sink term accounting e.g.
+for the evolution of ambient heat and moisture content, or sedimentation effects.
+Our approach also accounts for turbulent effects.
+Clouds in turbulent environments exhibit a complex interplay between microphysical processes and turbulent dynamics.
+Turbulence strongly influences the life cycle of clouds, as demonstrated in [Sha03].
+Turbulent models are essential for understanding droplet size spectra broadening [SPBC15, GA17], the mixing of cumulus clouds with their surroundings [AGP18, EKA+20], and the parameterization of mixed-phase clouds in global circulation models [FHFK14, FFB+16].
+In warm clouds, turbulence-induced localized temperature gradients and small-scale fluctuations can lead to supersaturation variations.
+These variations can activate haze particles even in environments with a mean subsaturation, and similarly, cloud particles may deactivate in saturated conditions if the fluctuations are sufficiently strong [PSK+20, AGP18].
+Therefore, microphysical fluctuations are crucial for accurately representing droplet size spectra in turbulent cloud environments.
+We represent these effects by means of stochastic parameterizations involving Brownian motions [CSG11, LC23] and possibly depending on the particle’s size.
+Thus, this research explores the activation and deactivation of cloud droplets in warm clouds, offering new theoretical perspectives.
+We aim to find analytical formulas for describing the cloud’s DSDs when both small cloud droplets and haze particles coexist, assuming condensation as the only growth mechanism.
+To this end, we add a sink term to the classical formulation of Köhler’s equation in order to parameterize the interaction of a population of droplets or other size-depleting mechanisms, like water vapor consumption or sedimentation [AS17, HKAA15, Kru20].
+The resulting equation describing condensational growth of droplets composing a monodisperse aerosol population allows for multiple stable states to co-exist besides the “haze" state, meaning the system exhibits multistability.
+Although the haze state remains stable, this multistability suggests that activated droplets can get stuck around a specific size due to the balancing act between two factors: the Köhler instability and mechanisms that decrease supersaturation or remove particles; see also [KM03, AS17, CSY+20].
+This is where turbulence effects throw a wrench in the picture.
 The stable states become “fuzzy states” that are encountered statistically (metastable) at various rates of occurrence.
 This means that particles can switch between activated and deactivated states, similar to how Brownian particles move around in a bumpy landscape.
-These turbulent jolts also significantly impact the activation-deactivation process into the form of various possible hysteresis loops (see Fig.
-5 below).
+These turbulent jolts also strongly impact the activation-deactivation process into the form of various possible hysteresis loops; see Section "Hysteresis effects" below.
 As a result, droplets can become activated at lower levels of supersaturation compared to the standard Köhler threshold.
-Finally, our framework allows us to derive analytical formulas for the size distribution of droplets (DSDs) within a cloud volume, accounting for these stochastic turbulent effects.
-In particular, our stochastic model predicts that droplet size distributions in Pi cloud convection chamber experiments behave like Gibbs states.
-It is shown that this prediction closely matches the actual DSDs observed in these experiments (see Fig.
-6 below).
+Finally, our framework allows us to derive analytical formulas for the DSDs within a cloud volume, accounting for these stochastic turbulent effects.
+In particular, our stochastic model predicts that the DSDs in the Pi cloud convection chamber experiments behave like Gibbs states, i.e.
+an equilibrium probability distribution that maximizes the system’s entropy subject to constraints on its average energy like in statistical mechanics [JKO98].
+It is shown that our theoretical predictions based on this formalism closely match the actual DSDs observed in these experiments; see Section "Pi-Chamber empirical distributions vs Gibbs states" below.
 Results
@@
 The theory of Köhler is the thermodynamical backbone to explain the formation of cloud droplets out of suspended aerosols in the atmosphere.
-It establishes the critical relative humidity threshold beyond which suspended particles activate and grow by diffusion (condensation) Köhler (1936).
+It establishes the critical relative humidity threshold beyond which suspended particles activate and grow by diffusion (condensation) [Köh36].
 The critical threshold is above saturation (hence, the supersaturation term is used), and it depends on the initial size and chemical composition of the particle in question.
-These particles are, according to Köhler, present in two distinct/dichotomical thermodynamic states: The first is known as haze and it is observed when the aerosol is humidified, although remains at equilibrium with its moist environment.
+These particles are, according to Köhler, present in two distinct/dichotomical thermodynamic states: The first is known as haze and it is observed when the aerosol is humidified, although it remains at equilibrium with its moist environment.
 The second state is referred to as activated droplet, whereby the particle harvests the available moisture by diffusion and, in principle, grows indefinitely.
-However, in nature, the available water molecules that are in supersaturation are limited, and in many cases, droplets and haze coexist in the cloud, and particles can oscillate between the two states.
-Particularly, in cases of small, warm clouds Hirsch et al.
-(2017); Altaratz et al.
-(2021).
-The critical relative humidity threshold needed for activation of a given aerosol with a solubility constant $k$ and a dry radius $r_{d}$ is analytically obtained using the Köhler’s equation Köhler (1936); Rogers and Yau (1989).
+However, in nature, the available water vapor molecules that are in supersaturation are limited, and in many cases, droplets and haze coexist in the cloud, and particles can oscillate between the two states.
+Particularly, in cases of small, warm clouds [HKA+17, AKA+21].
+The critical relative humidity threshold needed for activation of a given aerosol with a solubility constant $k$ and a dry radius $r_{d}$ is analytically obtained using Köhler’s equation [Köh36, RY89].
 Here, to simplify some derivations, $\lambda$ denotes the part of the relative humidity above $100\%$.
-If $r(t)$ denotes the radius of a given droplet at time $t$, its evolution is approximated by Köhler (1936); Arabas and Shima (2017):
+If $r(t)$ denotes the radius of a given droplet at time $t$, its evolution is approximated by [Köh36, AS17]:
 $\begin{cases}\dot{r}=\frac{D}{r}(\lambda-\lambda_{eq}(r))\\ \lambda_{eq}(r)=\frac{A}{r}-\frac{B}{r^{3}},\end{cases}$ (1)
-where $D$ is the diffusivity constant, $A$ relates to the water surface tension, $B=kr_{d}^{3}$, $\lambda$ is the ambient supersaturation and $\lambda_{eq}(r)$ determines the equilibrium supersaturation at the surface of a droplet of radius $r$; commonly known as Köhler curve Wallace and Hobbs (2006).
+where $D$ is the diffusivity constant, $A$ relates to the water surface tension, $B=kr_{d}^{3}$, $\lambda$ is the ambient supersaturation and $\lambda_{eq}(r)$ determines the equilibrium supersaturation at the surface of a droplet of radius $r$; commonly known as Köhler curve [WH06].
 Equation (1) shows that the particle is in equilibrium when $\lambda=\lambda_{eq}(r)$.
@@
 $\begin{cases}\dot{X}=\lambda-f(X),\\ f(X)=A(2DX)^{-1/2}-B(2DX)^{-3/2},\end{cases}$ (2)
-with $f(X)$ denoting the equilibrium supersaturation $\lambda_{eq}(r)$ as function of $X$.
-A rudimentary analysis of $f(X)$ reveals a single global maximum located at $X_{K}=3B/(2DA)$, and with critical supersaturation value $\lambda_{K}=f(X_{K})=(4A^{3}/27B)^{1/2}$, see also (Wallace and Hobbs, 2006, §6.1).
+with $f(X)$ denoting the equilibrium supersaturation $\lambda_{eq}(r)$ as function of (the scaled squared radius) $X$.
+A rudimentary analysis of $f(X)$ reveals a single global maximum located at $X_{K}=3B/(2DA)$, and with critical supersaturation value $\lambda_{K}=f(X_{K})=(4A^{3}/27B)^{1/2}$, see also Chapter 6.1 in [WH06].
 An example of Köhler curve is shown in Fig.
@@
 This curve reaches its maximum at $X_{K}=6\cdot 10^{-3}\mathrm{s}$, or at $r^{2}=0.48\mathrm{\mu m}^{2}$.
-Table 1: Köhler curve’s parameters for NaCl.
+Table 1. Köhler curve’s parameters for NaCl.
 $k$ $r_{d}$ $(\mu\mathrm{m})$ $A$ $(\mu\mathrm{m})$ $D$ ($\mu\mathrm{m}^{2}s^{-1}$)
 $\mathrm{NaCl}$ $1.28$ $5\times 10^{-2}$ $10^{-3}$ $40$
-For a given dry radius $r_{d}$, when the supersaturation $\lambda$ exceeds the corresponding $\lambda_{K}$, the particle is activated into a cloud droplet and grows by diffusion.
-On the contrary, for a given ambient supersaturation $\lambda$, particles whose radii are less than $(2DX_{u})^{1/2}$ remain stable in a haze state; see Fig.
-1a).
+For a specified dry radius, $r_{d}$, a particle will activate into a cloud droplet when the ambient supersaturation, $\lambda$, surpasses its critical supersaturation, $\lambda_{K}$.
+This activation process occurs through diffusional growth.
+Conversely, for a given supersaturation less than $\lambda_{K}$, particles with radii smaller than $(2DX_{u})^{1/2}$ will remain in a stable haze state.
+These equilibrium points are visually represented by the intersection of a horizontal line at the level of $\lambda$ and the Köhler curve, as depicted in Figure 1a.
+The solid and empty dots within this figure denote the stable and unstable equilibria, respectively.
 While Köhler theory is established for single particles, clouds consist of large families of particles coexisting in the same humidity environment.
-When a collection of droplets consumes the available supersaturation, the particles may cease growing by condensation, and the droplet size distributions (DSDs) stagnate and stop being displaced towards larger sizes.
+When a collection of droplets consumes the available supersaturation, the particles may cease growing by condensation, and the DSDs stagnate and stop being displaced towards larger sizes [KM03, SBK17].
 Here we modify the Köhler equation to the case of many particles in conditions of small warm clouds, when the available supersaturation is a limit factor.
 The proposed model extends the condensational growth equation—governed by the Köhler curve— by coupling all particles to supersaturation consumption and by adding stochastic disturbances, to account for (micro) small-scale turbulent effects.
-Figure 1: Panel a): Köhler curve associated with $\mathrm{NaCl}$; see Table 1 for the parameter values.
+Figure 1.
+Köhler and M-Köhler curves, and underlying potential.
+Panel a): Köhler curve associated with $\mathrm{NaCl}$; see Table 1 for the parameter values.
 A dry radius of $50\mathrm{nm}$ is prescribed.
@@
 For the cyan curve are shown furthermore its intersection with a given supersaturation $\lambda$ in between $\lambda_{c}$ and $\lambda_{h}$.
-Each filled dot corresponds to a stable equilibrium while the the empty dot corresponds to the unstable equilibrium.
+Each filled dot corresponds to a stable equilibrium while the empty dot corresponds to the unstable equilibrium.
 Panel c): The potential function associated with the cyan M-Köhler curve for the supersaturation value as indicated in the title.
@@
 In cases with relatively small supersaturation, as particles grow, they may consume the available supersaturation and therefore reduce their growth rate until reaching stagnation.
-In the realm of single-particle models these effects can be modeled by the introduction of a sink function into Eq.
+In the realm of single-particle models, these effects can be modeled by the introduction of a sink function into Eq.
 (2):
@@
 The function $f-g$ is called a M-Köhler curve hereafter.
-The term $g$ should remain negative and should decrease as $X$ increases.
-The sink term in growth rate aims at (i) modeling the supersaturation budget, and (ii) parameterizing other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al.
-(2014); Altaratz et al.
+The term $g$ should remain negative and should decrease as $X$ increases, as its main role is for accounting for saturation mechanisms in the particle growth.
+More precisely, this sink term in growth rate aims to (i) model the supersaturation budget, and (ii) parameterize other stagnating mechanisms into the condensational growth rate, for instance, due to sedimentation [Kru20] or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth [HKL+14, AKA+21].
+The sink term can also relate to first-order fluctuations in supersaturation due to condensation.
+Indeed, when a sizable amount of particles coexist in the same cloud parcel, the constant supersaturation assumption has to be amended taking account changes in $\lambda$ in the course of time, assuming $\lambda$ to be time-dependent in the first equation of Eq.
 (2).
-The sink term can also relate to first-order fluctuations in supersaturation.
-Indeed, when a sizable amount of particles coexist in the same cloud parcel, the constant supersaturation assumption has to be amended taking account changes in $\lambda$ in the course of time, assuming $\lambda$ in Eq.
-(3) to be time-dependent.
 At a first order, we are concerned with describing the evolution of $\lambda(t)=\lambda+\lambda^{\prime}(t)$ with $\lambda^{\prime}(0)=0$, i.e.
 we are seeking for a model of the rate of change of the fluctuations $\lambda^{\prime}$ around a given value $\lambda$.
-Following Arabas and Shima (2017), the rate of change $\lambda^{\prime}$ can be approximated by the ratio $-\dot{\rho_{v}}/\rho_{vs}$, in which $\rho_{v}$ denotes the ambient vapour density away from the droplet while, $\rho_{vs}$, denotes the vapour density at saturation.
-Then, by expressing $\rho_{v}$ as $\rho_{v}=4\pi N\rho_{w}r^{3}/3$, accounting for the density of liquid water $\rho_{w}$ and the droplet volume concentration $N$ of identical (monodisperse) particles in the cloud parcel, we arrive at:
-$\dot{\lambda^{\prime}}=-\beta(N)\dot{\left(X^{\frac{3}{2}}\right)},$ (4)
+Following [AS17], the rate of change $\lambda^{\prime}$ can be approximated by the ratio $-\dot{\rho_{v}}/\rho_{vs}$, in which $\rho_{v}$ denotes the ambient vapor density away from the droplet while, $\rho_{vs}$, denotes the vapor density at saturation.
+Then, by expressing $\rho_{v}$ as $\rho_{v}=4\pi N\rho_{w}r^{3}/3$, accounting for the density of liquid water $\rho_{w}$ and the droplet number concentration $N$ of identical (monodisperse) particles in the cloud parcel, we arrive at:
+$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\lambda^{\prime}=-\beta(N)\frac{\,\mathrm{d}}{\,\mathrm{d}t}{\left(X^{\frac{3}{2}}\right)},$ (4)
 when written in the variable $X=r^{2}/(2D)$, where the constant $\beta(N)$ is given by
 $\beta(N)=\frac{2^{7/2}D^{3/2}\pi\rho_{w}}{3\rho_{vs}}N;$ (5)
-see (Arabas and Shima, 2017, Eqns. (9) and (10)).
-The integration of Eq. (4) yields to $\lambda(t)=\lambda-\beta(N)X^{3/2}$ which eventually leads to the following condensational growth model:
-$\dot{X}=\lambda-f(X)-\beta(N)X^{3/2}.$ (6)
+see Eqns. (9) and (10) in [AS17].
+By integrating each side of Eq.
+(4), we obtain $\lambda^{\prime}(t)=-\beta(N)X^{3/2}$.
+When this expression of $\lambda^{\prime}$ is used in $\lambda(t)=\lambda+\lambda^{\prime}(t)$ we arrive at $\lambda(t)=\lambda-\beta(N)X^{3/2}$ which leads, from the first equation in Eq.
+(2), to the following condensational growth model:
+$\dot{X}=\lambda(t)-f(X)=\lambda-f(X)-\beta(N)X^{3/2}.$ (6)
 Thus, the term $-\beta(N)X^{3/2}$ acts as the sink function $g$ in Eq.
@@
 (7) to have two positive solutions are easily derivable.
-Actually, a simple proposition inferring the existence of two positive steady states $X^{\ast}_{h}$ and $X^{\ast}_{c}$ (with $X^{\ast}_{c}\geq X^{\ast}_{h}$) to condensational growth model analogues to Eq.
-(6) for a sink function $-\beta(N)X^{\alpha}$ ($\alpha>1$) is derived in in Methods (see Proposition 1 therein).
+Actually, the existence of haze and activated steady states, $X^{\ast}_{h}$ and $X^{\ast}_{c}$ (with $X^{\ast}_{c}\geq X^{\ast}_{h}$), can be inferred for condensational growth models analogue to Eq.
+(6) with more general sink functions $-\beta(N)X^{\alpha}$ ($\alpha>1$); see Proposition 1 in Section "Multistability" in Material and Methods.
 In any event, as shown in Fig.
 1b) the locations of $X^{\ast}_{h}$ and $X^{\ast}_{c}$ depend on the strength of the sink function’s coefficient $\beta(N)$.
-We define then the following supersaturation values $\lambda_{h}=f(X^{\ast}_{h})-g(X^{\ast}_{h})$, which corresponds to the local maximum of $f-g$, and $\lambda_{c}=f(X^{\ast}_{c})-g(X^{\ast}_{c})$, corresponding to its local minimum.
+We define then the following supersaturation values $\lambda_{h}=f(X^{\ast}_{h})-g(X^{\ast}_{h})$, corresponding to the local maximum of $f-g$, and $\lambda_{c}=f(X^{\ast}_{c})-g(X^{\ast}_{c})$, corresponding to its local minimum.
 In particular, $\lambda_{h}\geq\lambda_{c}$.
-Due to the presence of this new local minimum (compared to the Köhler curve), as supersaturation $\lambda$ varies, the number of equilibria varies through two saddle-node bifurcations, as sumarized in Fig.
-2.
-Figure 2: Going through multiple equilibria as $\lambda$ varies.
+Due to the introduction of a new local minimum compared to the original Köhler curve (Fig.
+1a), the number of equilibria varies with changes in supersaturation, $\lambda$, through two saddle-node bifurcations.
+As depicted in Figure 2, these bifurcations occur when the haze equilibrium (solid blue dot) approaches and merges with the unstable node (empty circle).
+Conversely, as $\lambda$ decreases, a similar bifurcation takes place when the activated equilibrium approaches the unstable node.
+The solid and purple dots in the bottom panels denote the critical values of $\lambda$ at which these bifurcations occur.
+Figure 2.
+Going through multiple equilibria as $\lambda$ varies.
 Two saddle-node (SN) bifurcations occur, one at $\lambda=\lambda_{c}$ and one $\lambda=\lambda_{h}$.
-For $\lambda<\lambda_{c}$, there is only a hazy stable equilibrium $X_{h}$ (blue dot in the leftmost upper diagram).
+For $\lambda<\lambda_{c}$, there is only a haze stable equilibrium $X_{h}$ (blue dot in the leftmost upper diagram).
 As one crosses from below the SN bifurcation point at $\lambda=\lambda_{c}$, two other equilibria emerge for a total of 3 equilibria coexisting for $\lambda_{c}<\lambda<\lambda_{h}$ (middle upper diagram): an activated stable equilibrium $X_{c}$ (red dot), an unstable equilibrium $X_{u}$ (open circle), and a haze equilibrium (blue dot).
-At the SN bifurcation occurring at $\lambda=\lambda_{h}$, $X_{u}$ and $X_{h}$ collide (purple dot on the rightmost lower diagram), and only the activated equilibrium survives as $\lambda>\lambda_{h}$ (red dot in the rightmost upper diagram).
-These two saddle-node bifurcations provide the nonlinear backbone for hysteresis to take place as discussed in Hysteresis effects, below.
-These hysteresis effects are well-known to be revealed in presence of stochastic disturbances; see e.g.
-Berglund (2011).
+At the SN bifurcation occurring at $\lambda=\lambda_{h}$, $X_{u}$ and $X_{h}$ “collide” (purple dot on the rightmost lower diagram), and only the activated equilibrium survives as $\lambda>\lambda_{h}$ (red dot in the rightmost upper diagram).
+These two saddle-node bifurcations serve as the nonlinear foundation for the hysteresis phenomenon discussed in the following Section on "Hysteresis effects." These hysteresis effects are well-known to be revealed in presence of stochastic disturbances combined with slow parameter variations; see e.g.
+[Ber11].
 We discuss below the physical motivation of introducing such disturbances in Eq.
-(6) or more generally, Eq.
+(6) or more generally, in Eq.
 (3).
 Brownian models and turbulent effects
-In a developing cloud, turbulent updrafts and mixing with its neighboring air provoke supersaturation fluctuations that naturally affect droplet growth and, consequently, the collective droplet size spectra Grabowski and Abade (2017); Abade, Grabowski, and Pawlowska (2018).
-Stochastic models for the velocity fluctuations and the supersaturation fields have been considered in previous studies and have shown various degrees of relevance Sardina et al.
-(2015); Siewert, Bec, and Krstulovic (2017).
-For a cloud parcel subject to turbulent updrafts, the authors in Sardina et al.
-(2015) prove that activated droplet size variance increases, at short times, proportionally to the square-root of time, analogous to Einstein’s diffusion formula.
+In a developing cloud, turbulent updrafts and mixing with its neighboring air provoke supersaturation fluctuations that naturally affect droplet growth and, consequently, the collective droplet size spectra.
+In that respect, we refer to [DSS+12, SS17] for observational studies in stratocumulus clouds, and to [PS09, KRZC97] regarding the influence of temperature and pressure fluctuations in supersaturation distribution.
+Stochastic models for the velocity fluctuations and the supersaturation fields have been considered in previous studies and have shown various degrees of relevance [SPBC15, SBK17].
+For a cloud parcel subject to turbulent updrafts, the authors in [SPBC15] prove that activated droplet size variance increases, at short times, proportionally to the square-root of time, analogous to Einstein’s diffusion formula.
 Their analytical findings are contrasted with direct numerical simulations of turbulence obeying the incompressible Navier-Stokes equation.
-In spite of this diffusion result with diverging variance, in Siewert, Bec, and Krstulovic (2017), the authors find a statistical steady state for droplet size, with DSDs characterized by exponential tails; see their Eq.
+In spite of this diffusion result with diverging variance, in [SBK17], the authors find a statistical steady state for droplet size, with DSDs characterized by exponential tails; see their Eq.
 (4.3).
@@
 Supersaturation being the driving parameter, we assume that it evolves around a given value $\lambda$ with fluctuations encoded by a stochastic term possibly depending on the particle size.
-For the sake of generality, we aim at modeling the inhomogeneous effects of supersaturation fluctuations on droplets of different size.
-Indeed, this size-dependent fluctuations appear, for instance, when taking the adiabatic limit of fast supersaturation fluctuations, as demonstrated in (Siewert, Bec, and Krstulovic, 2017, Eq.
-(4.3)).
-Hence, incorporating different supersaturation sources and sinks alter the expression for the noise law affecting droplet growth.
-Such noise inhomogeneity is here motivated from the fact that small fluctuations in droplet radius incur greater fluctuations in supersaturation when the particles are bigger.
-Contrary to haze particles, which are at a stable thermodynamic equilibrium and variations in their radius provoke smaller alterations in the supersaturation budget.
+The rationale behind this modeling assumption is that larger droplets are expected to have a greater impact on the supersaturation budget than the smaller ones, in a limited water vapor field.
+This is what we here referred to as inhomogeneous fluctuations, the inhomogeneous character being dependent on the droplets’ size.
+The type of inhomogeneity referred to here should not be confused thus with that caused by inhomogeneous mixing in clouds, occurring under large entrainment events [KKPF16].
 These considerations invite us to parameterize the fluctuations of $\lambda$ by means of a size-dependent stochastic term, $\sigma(X)\dot{W}_{t}$, leading to the following stochastic model:
 $\dot{X}=\underbrace{\lambda-f(X)+g(X)}_{\textstyle{-V_{\lambda}^{\prime}(X)}}+\sigma(X)\dot{W}_{t},$ (8)
-where $W_{t}$ is a Brownian motion (and $\dot{W}_{t}$ is Gaussian and white).
+where $W_{t}$ is a Brownian motion (and $\dot{W}_{t}$ is Gaussian and white). An example of the $\sigma(X)$ function is given in Section "Pi-Chamber empirical distributions vs Gibbs states" below.
 The function $V_{\lambda}$ is the potential function that collects the nonlinear, deterministic effects in the model.
@@
 (8).
-The equilibria of Eq.
-(3) correspond exactly to the local critical points of the potential function—see Fig.
+The equilibria of the deterministic counterpart of Eq.
+(8) (i.e.
+Eq.
+(3)) correspond exactly to the local critical points of the potential function—local minima in Fig.
 1c)—and their stability is determined by the local curvature of $V_{\lambda}$ at those points, see e.g.
-Berglund (2011); Pavliotis (2014).
+[Ber11, Pav14].
 Note that Eq.
-(8) can always be recast into a form in which the stochastic disturbance term becomes $X$-independent, at the expense of changing the potential function via the Lamperti transformation (Pavliotis, 2014, §3.6).
-This trick is recalled in Methods as it is also part of our model’s analysis below.
+(8) can always be recast into a form in which the stochastic disturbance term becomes $X$-independent, at the expense of changing the potential function via the Lamperti transformation; see Chapter 3.6 in [Pav14].
+This trick is recalled in Section "Lamperti transformation" in Material and Methods as it is also part of our model’s analysis below.
 For a family of monodisperse droplets in the same cloud volume, each one experiences different supersaturation fluctuations, albeit with the same mean and variance, as encapsulated by the term $\sigma(X)\dot{W}_{t}$ in Eq.
@@
 (8) provides the DSD at a given time $t$; see e.g.
-Pavliotis (2014); Risken (1989).
-The stationary solution to the FPE is then given analytically and known as the Gibbs state (Pavliotis, 2014, Eq.
-(4.35)):
+[Pav14, Ris89].
+The stationary solution to the FPE is then given analytically and known as the Gibbs state (see [JKO98] and Eq.
+(4.35) in [Pav14]):
 $\rho_{\lambda}(X)=Z_{\lambda}^{-1}\frac{\exp\bigg{(}2\int^{X}\frac{\lambda-f(x)+g(x)}{\sigma^{2}(x)}\,\mathrm{d}x\bigg{)}}{\sigma^{2}(X)},$ (9)
 where $Z_{\lambda}$ is a normalizing constant.
-This distribution is a generalization of what is found in (McGraw and Liu, 2006, Eq.
-(6)).
-Such equation yields a Weibull distribution for droplet radius $r$, without considering curvature and chemistry effects, i.e.
-Köhler theory, sink functions or state-dependent noise.
-One can, thus, derive the Weibull distribution $\rho_{W}$ for the radius $r$ by setting $f=g\equiv 0$ and $\sigma$ constant, where the resulting DSD is integrable only if there is mean subsaturation, $\lambda<0$:
+This distribution is a generalization of Weibull distribution found in Eq.
+(6) in [ML06] to explain diffusive broadening of the DSDs in turbulent clouds.
+Within our framework, it corresponds to Gibbs states without sink function and state-dependency in the noise, i.e.
+density distribution $\rho_{W}$ (for the radius $r$) obtained by setting $f=g\equiv 0$ and $\sigma$ constant.
+In this particular case, the resulting DSD is integrable only if there is mean subsaturation, $\lambda<0$ and is given by:
 $\rho_{W}(r)=Z_{\lambda}^{-1}re^{\gamma r^{2}},$ (10)
-where $\gamma$ is a parameter proportional to $\lambda$, droplet number concentration and liquid water fraction; see (McGraw and Liu, 2006, Eq.
-(8)).
+where $\gamma$ is a parameter proportional to $\lambda$, droplet number concentration and liquid water fraction; see Eq.
+(8) in [ML06].
 In comparison, $\rho_{\lambda}$ in Eq.
-(9) can be integrable if $\lambda$ is positive, provided that the sink function $g$ has suitable decay properties detailed in Proposition 2 found in The confining potential in Methods.
-For instance considering sedimentation, or any size-independent removal mechanism, would provide an integrable DSD; see Krueger (2020) and (Chandrakar et al., 2020, Sections 2.1, 2.2 and 2.3), where the resulting DSDs are an explicit evaluation of $\rho_{\lambda}$ in Eq.
-(9).
-For a broad class of $f$, $g$ and $\sigma$, the Gibbs state $\rho_{\lambda}$ exhibits multimodality intimately related to shape of the potential $V_{\lambda}$.
+(9) can be integrable if $\lambda$ is positive and subject to size-dependent fluctuations, provided that the sink function $g$ has suitable decay properties such as detailed in Proposition 2; see Section "The confining potential" in Material and Methods.
+We refer to [CSY+20] for other theoretical approaches to infer integrable DSDs for fluctuating supersaturation with size-dependent removal terms.
+Our framework, for a broad class of functions $f$, $g$ and $\sigma$, allows for Gibbs state $\rho_{\lambda}$ that exhibits multimodality intimately related to shape of the potential $V_{\lambda}$.
 More precisely, the Gibbs state’s modes, the local maxima of $\rho_{\lambda}$, are found by solving the following equation:
 $-V^{\prime}_{\lambda}(X)=\frac{1}{2}\sigma^{\prime}(X)\sigma(X).$ (11)
-See Noise-induced metastability in Methods for the derivation of this equation.
+See Section "Noise-induced metastability" in Material and Methods for the derivation of this equation.
 This fundamental equation teaches us how the Gibbs state modes result from the interaction between the noise term and the potential function.
 The Gibbs state’s modes correspond to metastable states for which the random fluctuation term, $\sigma(X)\dot{W}_{t}$, in Eq.
-(8) acts as a source for the particles to experience transitions from a statistically typical droplet size to another one; see schematic of Fig.
-3.
+(8) acts as a source for the particles to experience transitions from a statistically typical droplet size to another one.
 When the noise is state-independent, i.e.
-when $\sigma(X)=\sqrt{2\varepsilon}$, then $\sigma^{\prime}(X)=0$ and the metastable states are directly determined by the zeros of $-V_{\lambda}^{\prime}$.
+when $\sigma(X)=\sqrt{2\varepsilon}$, for some positive constant $\varepsilon>0$ (controlling the noise variance), then $\sigma^{\prime}(X)=0$ and the metastable states are directly determined by the zeros of $-V_{\lambda}^{\prime}$.
 In this case, for Eq.
 (6), the Gibbs states are bimodal when $\lambda_{c}<\lambda<\lambda_{h}$, and unimodal when there is just one underlying stable equilibrium, i.e.
-for either $\lambda<\lambda_{c}$ or $\lambda>\lambda_{h}$; see Fig.
+for either $\lambda<\lambda_{c}$ or $\lambda>\lambda_{h}$, as shown in Fig.
+2, whereby the solid points denote the presence of equilibria, ranging from a single haze equilibrium to an activated one.
+When $\lambda_{c}<\lambda<\lambda_{h}$, there is a notable chance of encountering haze or activated droplets.
+As depicted in Fig.
+3, the red curve represents a potential energy landscape with two local minima similar to that shown in Fig.
+1c).
+Stochastic fluctuations can then overcome the potential barrier between these two minima, leading to activation even when average supersaturation falls below the Köhler threshold.
+This phenomenon aligns with real-world scenarios where turbulent supersaturation variations can induce haze activation [PSK+20].
+Figure 3.
+Cartoon of haze-to-droplet stochastic transitions.
+Here $\lambda$ is kept fixed and the situation depicted corresponds to $\lambda$ in $(\lambda_{c},\lambda_{h})$ where two stable states coexist (haze particle and activated droplet); see Fig.
 2.
-When $\lambda_{c}<\lambda<\lambda_{h}$, there is a non-vanishing probability of encountering haze or activated droplets; see Fig.
-3.
-This property is consistent with realistic situations in which turbulent supersaturation fluctuations make haze susceptible of activating even if on average the Köhler’s threshold is not attained Prabhakaran et al.
-(2020).
-Figure 3: Cartoon of haze-to-droplet stochastic transitions.
-Here $\lambda$ is kept fixed and the situation depicted corresponds to $\lambda$ in $(\lambda_{c},\lambda_{h})$ where coexist two stable states (haze particle and activated droplet); see Fig.
-2.
-Our modeling allows for other useful analytic insights.
+Our modeling framework allows for other useful analytic insights.
 For instance, the expected residence times at the haze or activated state can be precisely determined.
 They are indeed known to relate to the potential barrier, $V_{\lambda}(X_{u})-V_{\lambda}(X_{h})$ (see Fig.
-1c), according to the Kramers’ time formulas Kramers (1940); Berglund (2011).
-The mean activation time $\tau_{act}$ is then given by (Matkowsky and Schuss, 1981, Eq.
-(2.19))
+1c), according to the Kramers’ time formulas [Kra40, Ber11].
+The mean activation time $\tau_{act}$ is then given by Eq.
+(2.19) in [MS81]
 $\tau_{act}(\lambda)=\frac{2\pi}{\sqrt{|V_{\lambda}^{\prime\prime}(X_{u})|V_{\lambda}^{\prime\prime}(X_{h})}}e^{\frac{V_{\lambda}(X_{u})-V_{\lambda}(X_{h})}{\varepsilon}}.$ (12)
-Therefore, the closer $\lambda$ is to the critical value $\lambda_{h}$ from below, the closer $X_{u}$ is to a collision with $X_{h}$ (see Fig.
+Therefore, the closer $\lambda$ is to the critical value $\lambda_{h}$ from below, the closer $X_{u}$ is to $X_{h}$ (see Fig.
 2) and the shallower the potential barrier becomes, resulting into more common particle activation.
@@
 Hysteresis effects
-Figure 4: Rationale of an hysteresis path.
+Figure 4.
+Rationale of an hysteresis path.
 As $\lambda$ varies, the potential landscape changes and the stochastic disturbances can help trigger the critical transition as one approaches the tipping points, here the saddle-node bifurcation points.
-Figure 5: Panels a), b) and c): The thin gray line indicates the hysteresis loops for different rates of change of supersaturation $\dot{\lambda}$ as indicated in the title.
+Figure 5.
+A landscape of hysteresis loops.
+Panels a), b) and c): The thin gray line indicates the hysteresis loops for different rates of change of supersaturation $\dot{\lambda}$ as indicated in the title.
 The noise variance is constant for these three panels as indicated in the legend of Panel a).
@@
 For all panels, the cyan line is the M-Köhler curve associated with Eq.
-(6) and $\beta(N)=3.6\times 10^{-2}\mathrm{s}^{-3/2}$.
-The hysteresis behavior pointed out at the end of Multistable Köhler curves is now analyzed when the supersaturation parameter $\lambda$ is allowed to drift through the critical values $\lambda_{c}$ and $\lambda_{h}$, slower compared to the stochastic fluctuations.
+(6) where $\beta(N)=3.6\times 10^{-2}\mathrm{s}^{-3/2}$.
+The hysteresis behavior pointed out at the end of Section "Multistable Köhler curves" is now analyzed when the supersaturation parameter $\lambda$ is allowed to drift through the critical values $\lambda_{c}$ and $\lambda_{h}$, slower compared to the stochastic fluctuations.
 Such relatively slower variations in $\lambda$ are associated with buoyancy-driven supersaturation changes.
@@
 Indeed, when a haze particle undergoes activation— starting at point A and following the red, wobbly line—, its deactivation will occur at much lower supersaturation levels— as seen when following the blue wobbly line that ends in point B—.
-Similar cycles have also been reported in the absence of aerosol curvature and chemistry effects, and uniquely due to noisy supersaturation fluctuations around a subsaturated mean; see (Siewert, Bec, and Krstulovic, 2017, Fig.
-4).
+Similar cycles have also been reported in the absence of aerosol curvature and chemistry effects, and uniquely due to noisy supersaturation fluctuations around a subsaturated mean; see e.g.
+Fig.
+4 in [SBK17].
 Within the present framework, it remains to investigate the effects of (i) different rates of changes in supersaturation and (ii) variations in the magnitude of the fast supersaturation fluctuations as encapsulated by the noise term.
-The hysteresis phenomenon under the presence of noise has been investigated in previous work Berglund and Gentz (2002), where it is found that a large (additive) noise variance $\sigma$, on average, reduces the area of hysteresis loops.
+The hysteresis phenomenon under the presence of noise has been investigated in previous work [BG02], where it is found that a large (additive) noise variance $\sigma$, on average, reduces the area of hysteresis loops.
 In our context, this implies that droplet activation can occur at much lower supersaturation values compared to $\lambda_{h}$.
-To numerically examine these ideas, we solve the Eq.
+To numerically examine these ideas, we solve Eq.
 (8) associated with the sink term $g(X)=-\beta(N)X^{3/2}$ used in Eq.
-(6), for specific noise variances and mean supersaturation $\lambda$ ranging from $0.06\%$ to $0.11\%$ with different linear rate $\dot{\lambda}$.
-The numerical experiments regarding hysteresis are shown in Fig.
-5, where Eq.
-(6) was integrated using an Euler-Maruyama scheme with a time step of $10^{-2}\mathrm{s}$.
-In Fig.
-5a), b) and c), are shown hysteresis paths solving Eq.
-(6) for three rates of change $\dot{\lambda}$.
+(6), for specific noise variances and mean supersaturation $\lambda$ ranging from $0.06\%$ to $0.11\%$ with different rate of change $\dot{\lambda}$.
+We refer to Section "Hysteresis path algorithm" in Material and Methods for numerical details.
+Our model supports different hysteresis cycles as shown in Figure 5.
+The nature of these hysteresis loops is controlled by the rate of change of supersaturation $\dot{\lambda}$ and the intensity of the turbulent effects.
+Figure 5a) to Figure 5c) correspond to an increasingly faster response of droplet growth to supersaturation changes.
+Figure 5d) to Figure 5f) correspond to an increasing susceptibility to activation as the noise variance is increased, i.e.
+the turbulent effects are more pronounced while the rate of change $\dot{\lambda}$ is kept fixed.
 As expected, when the rate of change in supersaturation is not slow enough, it can prevent droplet activation— see Fig.
@@
 5b) and c), we observe that only when the cooling rate is slowest do we observe a droplet growth attached to the Köhler curve.
-Panels d), e) and f) of Figure 5 provide an illustration of results of Berglund and Gentz (2002) for Eq.
-(6).
-Here we show the effects of increasing turbulent noise intensity as a haze particle grows and activates in a slowly cooling air volume.
-In such case we are in the situation of Fig.
-5c) where the particle’s size is dictated by the Köhler curve— modulo zero-mean fluctuations—.
-As noise variance is increased, the hysteresis loop is collapsed to the extent that droplet activation is expected as soon as $\lambda$ surpasses $\lambda_{c}$.
-Consequently, it is expected that slowly lifted humid parcels undergoing heaving turbulent updrafts— here encoded in the noise variance— can contain particles that activate at much lower supersaturations compared to that predicted by Köhler.
+Panels d), e) and f) of Figure 5 provide for Eq.
+(6), illustrations of hysteresis cycles that are consistent with those analyzed for large-noise regimes in theoretical studies [BG02].
+Compared to Panels a) and b) of Figure 5, the rate of change of supersaturation $\dot{\lambda}$ is smaller, resulting into sharper transitions to activation similar than those shown in Figure 5c).
+This combined with a larger turbulent noise intensity than in the cases of Figure 5a)-b)-c), may provoke transitions to occur before reaching the haze-to-droplet saddle-node bifurcation, i.e.
+particles that activate at much lower supersaturations compared to that predicted by Köhler.
+Physically it corresponds to haze particles that grow in a slowly cooling air volume embedded in a highly turbulent environment undergoing heaving turbulent updrafts.
+In some dramatic limit, the hysteresis loop may collapse such as shown in Figure 5f).
+This extreme situation is to put in contrast with the other extreme situation shown in Figure 5a) where activation is "unrealized" due to faster rate of change of supersaturation and weakly turbulent updraft.
 Pi-Chamber empirical distributions vs Gibbs states
-Small, warm convective clouds are abundant in the atmosphere but hard to measure Koren et al.
-(2008).
-Such clouds are often characterized by weak updrafts and small supersaturation, allowing for a significant part of the aerosol to coexist as haze together with the activated part Altaratz et al.
-(2021); Hirsch et al.
-(2017).
-Cloud chambers are often used to study clouds in such thermodynamic conditions experimentally Prabhakaran et al.
-(2020); Chandrakar et al.
-(2016).
+Small, warm convective clouds are abundant in the atmosphere but hard to measure [KOF+08].
+Such clouds are often characterized by weak updrafts and small supersaturation, allowing for a substantial part of the aerosol to coexist as haze together with the activated part [AKA+21, HKA+17].
+Cloud chambers are often used to study clouds in such thermodynamic conditions experimentally [PSK+20, CCC+16].
 In such work, the so-called Pi-chamber is seeded with monodisperse families of particles, and turbulent fluctuations are induced by changes in temperature, pressure, and aerosol injection rates.
-Disregarding transients, the system reaches a steady state in number concentration and size distributions are calculated spanning hazy and active particles.
-Depending on the supersaturation mean and variance, DSDs display different characteristics involving their multimodality and spread; see (Prabhakaran et al., 2020, Fig.
-2).
+Disregarding transients, the system reaches a steady state in number concentration and DSDs are calculated spanning haze and active particles.
+Depending on the supersaturation mean and variance, DSDs display different characteristics involving their multimodality and spread; see Fig.
+2 in [PSK+20].
 We show here that such experimental distributions are well approximated by the Gibbs states of our Brownian modelling framework once the sink and noise terms are appropriately designed.
-Our goal is to provide a qualitative and quantitative correspondence between the present theory and the experiments from Prabhakaran et al.
-(2020).
+Our goal is to provide a qualitative and quantitative correspondence between the present theory and the experiments from [PSK+20].
 We first address the sink term design.
-Note that for a monodisperse family of particles embedded in a cloud parcel, the supersaturation budget must be included in the collective droplet growth Squires (1956); Korolev and Mazin (2003).
-Because the Pi-chamber is closed and in steady-state, the sink term affecting supersaturation is determined by water vapor condensation onto the particles.
-The mean sources of supersaturation are taken to be zero, apart from fluctuations encoded in $\sigma(X)$; see Eq.
-(8).
-Consequently, and to a first approximation, the initially prescribed supersaturation $\lambda$ is consumed proportionally to the radius of the droplets, so that it evolves according to:
-$\lambda(t)\approx\lambda-\beta X(t)^{1/2}+\sigma(X)\dot{W}_{t},$ (13)
-where the absorption parameter $\beta$ depends on the number of active droplets, and the effective diffusivity of water vapour in air Khain and Pinsky (2018).
-By substituting $\lambda$ in the (standard) Köhler equation (1), the new term $-\beta X(t)^{1/2}$ plays the role of a sink term in Eq.
-(3), so that $g(X)=-\beta X^{1/2}$.
-Regarding the term $\sigma(X)\dot{W}_{t}$, the inhomogeneous effect of fluctuations in haze and cloud droplets becomes apparent by qualitatively examining the DSDs obtained in the Pi-chamber; see (Prabhakaran et al., 2020, Fig.
-2).
-To account for such dependence on particle size, we propose the following stochastic parameterization:
+Note that for a monodisperse family of particles embedded in a cloud parcel, the supersaturation budget must be included in the collective droplet growth [Squ56, KM03].
+Because the Pi-chamber is in steady-state, the sink term affecting supersaturation is determined by water vapor condensation onto the particles, which boils down to assuming that supersaturation is consumed proportionally to the radius of the droplets.
+Thus a sink term of type $g(X)=-\beta X^{1/2}$ is chosen in Eq.
+(8), where the absorption parameter $\beta$ depends on the number of active droplets, and the effective diffusivity of water vapor in air [KP18].
+It must be noted, however, that although in steady-state, the Pi-chamber is not a closed system.
+There is an intrinsic decorrelation timescale of supersaturation; see [CCC+16].
+This timescale is usually determined by the coefficients in front of the deterministic terms of a (linear) stochastic Langevin equation governing the supersaturation evolution; see e.g.
+Eq.
+(2) in [CCC+16].
+In our formalism, while not solving an explicit governing equation for supersaturation we can still determine a decorrelation timescale for supersaturation albeit in an indirect fashion.
+One first solves Eq.
+(8), records $X(t)$ and then form supersation time series according to:
+$\lambda(t)=\lambda-\beta X(t)^{1/2}+\sigma(X(t))\dot{W}_{t}.$ (13)
+Now the expectation of the stochastic term $\sigma(X)\dot{W}_{t}$ is zero due to the martingale property of Brownian motion and the independence of $\sigma(X(t))$ and $\dot{W}_{t}$.
+This is actually a fundamental property of stochastic differential equations driven by Brownian motion; see Section 4.2.6 in [Gar09].
+Therefore the expected value of supersaturation has a decorrelation time given by that of $\lambda-\beta\mathbb{E}X(t)^{1/2}$, that is a function of the decorrelation time of the droplet radius.
+Thus the droplet size effects on supersaturation fluctuations encoded by the stochastic term $\sigma(X)\dot{W}_{t}$ affect the decorrelation time of the expected value of supersaturation in an indirect way, after solving Eq.
+(8) to get access to $X(t)$.
+In the case of the Pi-chamber datasets analyzed below, we model the size-dependent effects on supersaturation fluctuations via the following functional:
 $\sigma(X)=\Big{(}\sigma_{1}+\frac{\sigma_{2}-\sigma_{1}}{2}\big{(}1+\psi(X)\big{)}\Big{)},$ (14)
-$\psi(X)=\tanh(2\kappa D(X-X^{\ast})).$
-Here, the hyperbolic tangent is employed as a step function that allows for interpolating smoothly between the noise strengths that affect haze and activated particles, $\sigma_{1}$ and $\sigma_{2}$ respectively.
-Thus, we require that $\sigma_{2}>\sigma_{1}>0$ for the reason stated above.
-The value of the slope is chosen to be sufficiently large— i.e., $102D\kappa=$ 10— so that the hyperbolic tangent approximates a step function, while still exhibiting the attributes of a smooth function, more amenable for analysis and numerical simulations.
-The “ignition” parameter $X^{\ast}$ is chosen to be equal to $6.2\cdot 10^{-3}\mathrm{s}$ or, in diameter, $d^{\ast}=1.41\mu\mathrm{m}$, which yields a correct location of the haze modes, and it is in the order of the critical Köhler diameter which is $d_{K}=1.7\mu\mathrm{m}$.
-Such critical value is derived from Eq.
-(1), as explained in Köhler theory, for the curvature and solubility coefficients $A=1.4\cdot 10^{-3}\mu\mathrm{m}$ and $B=3.5\cdot 10^{-4}\mu\mathrm{m}^{3}$; see Prabhakaran et al.
-(2020).
+$\psi(X)=\tanh(2\kappa D(X-X^{\ast})),$
+where we require that $\sigma_{2}>\sigma_{1}>0$ so that larger droplets have a bigger impact on supersaturation fluctuations than the smaller ones.
+The value of the slope is chosen to be sufficiently large — i.e., $102\kappa D=$ 10 — so that the hyperbolic tangent function, $\psi$, approximates a step function, while still exhibiting the attributes of a smooth function, more amenable for analysis and numerical simulations.
+The “ignition” parameter $X^{\ast}$ is chosen to be equal to $6.2\cdot 10^{-3}\mathrm{s}$ or, in diameter, $d^{\ast}=1.41\mu\mathrm{m}$, which yields a correct location of the haze mode, and it is in the order of the critical Köhler diameter which is $d_{K}=1.7\mu\mathrm{m}$.
+Such a critical value is derived from Eq.
+(1), as explained in Köhler theory, for the curvature and solubility coefficients $A=1.4\cdot 10^{-3}\mu\mathrm{m}$ and $B=3.5\cdot 10^{-4}\mu\mathrm{m}^{3}$; see [PSK+20].
 We, thus, arrive at the following Brownian model to describe cloud droplet size evolution in the Pi-chamber:
@@
 Eq. (16) plays a key role to estimate $\beta$: once $\beta$ is known, the only parameters left for calibration of the model are $\sigma_{1}$, $\sigma_{2}$ and the stiffness parameter $\kappa$.
-Figure 6: Panels a), b) and c): Köhler and M-Köhler curves (dashed and solid curves, respectively) for each supersaturation value considered in PI-Chamber Empirical Distribution vs Gibbs States (Cases I, II, and III).
+Figure 6.
+Pi-chamber experimental distributions vs Gibbs states, and underlying M-Köhler curves.
+Panels a), b) and c): Köhler and M-Köhler curves (dashed and solid curves, respectively) for each supersaturation value considered in Section "Pi-Chamber Empirical Distribution vs Gibbs States" (Cases I, II, and III).
 The parameters of the Köhler curve are $A=1.4\cdot 10^{-3}\mathrm{\mu m}$ and $B=3.5\cdot 10^{-4}\mathrm{\mu m^{3}}$ and the other parameters of model Eq.
 (15) are given in Table 2.
-The black dot denotes the location of the metastable state induced by the state-dependent noise; see Noise-induced metastability in Methods.
-Panels d), e) and f): Experimental DSDs from Prabhakaran et al.
-(2020) vs Gibbs states from Eq.
+The black dot denotes the location of the metastable state induced by the state-dependent noise; see Section "Noise-induced metastability" in Material and Methods.
+Panels d), e) and f): Experimental DSDs from [PSK+20] vs Gibbs states from Eq.
 (9).
 Both the data histograms and Gibbs states are normalized according to their integral.
-The insets show the transformed potentials (due to Lamperti transformation) explaining the shapes of the Gibbs states; see The Lamperti transformation in Methods.
+The insets show the transformed potentials (due to Lamperti transformation) explaining the shapes of the Gibbs states; see Section "The Lamperti transformation" in Material and Methods.
 We now confront the ability of our Brownian model (15) to approximate experimental DSDs by the Gibbs states formula (9).
-Three experimental setups are considered in that respect, following those of the Pi-chamber experiments from Prabhakaran et al.
-(2020) corresponding to different supersaturation mean and fluctuation properties reported in such paper.
-These are referred to as according to Prabhakaran et al.
-(2020): a) mean-dominated, b) fluctuation influenced and c) fluctuation dominated acivation.
+Three experimental setups are considered in that respect, following those of the Pi-chamber experiments from [PSK+20] corresponding to different supersaturation mean and fluctuation properties reported in such paper.
+These are referred to as according to [PSK+20]: a) mean-dominated, b) fluctuation influenced and c) fluctuation dominated acivation.
 Within our modeling framework, these regimes can be organized as follows:
   1. Case I: $\lambda>\lambda_{K}\rightarrow$ mean-dominated activation
-  2. Case II: $\lambda_{c}<\lambda<\lambda_{K}\rightarrow$ fluctuation influenced activation
-  3. Case III: $\lambda<1\rightarrow$ fluctuation dominated activation
-The parameter $\lambda_{K}$ is the critical Köhler threshold, which is derived to be $\lambda\approx 1.00108$, for the type of aerosols considered in Prabhakaran et al.
-(2020).
-The steady-state supersaturation value in each case is also derived from Prabhakaran et al.
-(2020).
+  2. Case II: $\lambda_{c}<\lambda<\lambda_{K}\rightarrow$ fluctuation-influenced activation
+  3. Case III: $\lambda<1\rightarrow$ fluctuation-dominated activation
+The parameter $\lambda_{K}$ is the critical Köhler threshold, which is derived to be $\lambda\approx 1.00108$, for the type of aerosols considered in [PSK+20].
+The steady-state supersaturation value in each case is also derived from [PSK+20].
 Their theoretical calculations reveal indeed that Case I corresponds to $\lambda=1\%$.
@@
 In Case II, the modal activated diameter is $d_{c}=9.141\mu\mathrm{m}$, giving $X_{c}=0.2611\mathrm{s}$ and $\beta=1.4\cdot 10^{-3}\mathrm{s}^{-1/2}$.
-These modal values are obtained from the data of Prabhakaran et al.
-(2020).
-For Case III, though, since the system is subsaturated and haze particles are more numerous—see Table 1 of Prabhakaran et al.
-(2020)—the term $\beta$ is practically zero.
+These modal values are obtained from the data of [PSK+20].
+For Case III, though, since the system is subsaturated and haze particles are more numerous—see Table 1 of [PSK+20]—the term $\beta$ is practically zero.
 The corresponding original Köhler and M-Köhler curves are shown in Fig.
 6a) and b), as dashed and solid curves, respectively, and shown as functions of the diameter $d=2\sqrt{2DX}$.
-For Case III, because there is not sink term, only the original Köhler curve is shown.
+For Case III, because there is no sink term, only the original Köhler curve is shown in Fig.
+6c).
 Once the parameters for the relative disturbances, $\sigma_{1}$, $\sigma_{2}$ and $\kappa$ are chosen as listed in Table 2, the corresponding Gibbs states $\rho_{\lambda}$ given by Eq.
-(9) exhibit striking skills in reproducing the empirical DSDs in each of the three Pi-chamber experiments taken from Prabhakaran et al.
-(2020).
+(9) exhibit striking skills in reproducing the empirical DSDs in each of the three Pi-chamber experiments taken from [PSK+20].
 The results are shown in the Figs.
@@
 The shapes of the Gibbs states are highly correlated to those of the transformed potentials shown in the insets of Fig.
-6, obtained by the Lamperti transformation; see Methods.
+6, obtained by the Lamperti transformation; see Section "The Lamperti transformation" in Material and Methods.
 We observe in particular that the potential wells’ locations match the Gibbs states modes’ locations.
@@
 (15) can be interpreted as a Brownian particle evolving within that potential whose fluctuations are just driven by a white noise, independently of the size of the particle.
-As sanity check, one observes that the experimental DSDs’ modes are well predicted by Eq.
+The theory allows us to make striking predictions.
+For instance, the experimental DSDs’ modes are well predicted by Eq.
 (11), as shown in Fig.
-7 of Noise-induced metastability, especially in Cases I and II, where the classical Köhler theory is unable to predict an activated mode, since $\lambda>\lambda_{K}$.
+7 of Section "Noise-induced metastability" in Material and Methods, especially in Cases I and II, where the classical Köhler theory is unable to predict an activated mode, since $\lambda>\lambda_{K}$.
 Noteworthy is the creation of a metastable haze-state in Case I resulting from the interaction between the noise term $\sigma(X)\dot{W}_{t}$ and the nonlinear drift terms $-f(X)-\beta X^{1/2}$ as revealed by the transformed potential shown in the inset of Figs.
-6d); see Noise-induced metastability in Methods.
-Table 2: Supersaturation $\lambda$ and parameter values of the sink term and state-dependent noise $\sigma(X)$ in Eq. (LABEL:eta_term_chamber).
+6d); see Section "Noise-induced metastability" in Material and Methods.
+Table 2. Supersaturation values, sink term’s and state-dependent noise’s parameters.
 […]
@@
 Discussion
-Köhler theory stands in the foundation of cloud physics.
+Köhler theory is at the basis of cloud physics.
 For a given dry aerosol size and composition, it predicts the particle’s state, haze, or activated droplet as a function of the particle size and the supersaturation.
@@
 The curve left to the mode describes a stable fixed point, whereas any particle that resides on the curve right to the mode is on an unstable fixed point such that a perturbation to the right (or upward) will shift the particle to a continuous growth by diffusion and a perturbation to the left (or downward) will send the particle to the haze state.
-Warm and small clouds that are driven by weak temperature or humidity perturbations Eytan et al.
-(2020); Altaratz et al.
-(2021); Hirsch et al.
-(2017) are superabundant and have an important radiative effect.
+Warm and small clouds that are driven by weak temperature or humidity perturbations [EKA+20, AKA+21, HKA+17] are abundant and have an important radiative effect.
 However, due to their sizes, weak optical signature, and short lifetimes, these clouds are mostly overlooked, and their properties are poorly understood.
-Such clouds often exist near the transition between haze to cloud Koren et al.
-(2009) a regime that can be studied in cloud chambers Prabhakaran et al.
-(2020); Chandrakar et al.
-(2016).
-Here, we address these clouds by extending Köhler theory to population of droplets embedded in turbulent environments.
+Such clouds often exist near the transition between haze to cloud [KFJA09] a regime that can be studied in cloud chambers [PSK+20, CCC+16].
+Here, we address these clouds by extending Köhler theory to a population of droplets embedded in turbulent environments.
 While activation is explained by Köhler’s theory, the reverse process, i.e.
 deactivation, is dynamically different.
-It has been documented in previous works Arabas and Shima (2017); Korolev and Mazin (2003) and in the present study that deactivation is intimately tied to the hysteresis phenomenon through which deactivation occurs at a smaller supersaturation compared to the critical Köhler value $\lambda_{K}$.
+It has been documented in previous works [AS17, KM03] and in the present study that deactivation is intimately tied to the hysteresis phenomenon through which deactivation occurs at a smaller supersaturation compared to the critical Köhler value $\lambda_{K}$.
 The addition of a sink term in the condensational growth equation — see Eq.
-(3)— yields hysteresis directly from an analytical perspective: in order to deactivate a family of monodisperse cloud droplets, the supersaturation needs to be brought strictly below the critical value $\lambda_{K}$ predicted by Köhler, see Figs.
-1b).
-The origin of this behavior is the presence of another saddle-node bifurcation that results from the bending of the M-Köhler curve introduced in Multistable Köhler curves This bending translates further into the appearance of another stable equilibrium on the activated droplet spectrum as a balance between condensational instability and sink factors, like supersaturation consumption or particle removal.
-We refer to the section Multistability in Methods for conditions on the sink function $g$ ensuring for such a bending to occur, and thus the creation of an extra saddle-node bifurcation in Eq.
-(3).
+(3) — yields hysteresis directly from an analytical perspective: in order to deactivate a family of monodisperse cloud droplets, the supersaturation needs to be brought strictly below the critical value $\lambda_{K}$ predicted by Köhler.
+The origin of this behavior is the presence of another saddle-node bifurcation that results from the bending of the M-Köhler curve introduced in Section "Multistable Köhler curves." This bending translates further into the appearance of another stable equilibrium on the activated droplet spectrum as a balance between condensational instability and sink factors, like supersaturation consumption or particle removal.
 The introduction of Brownian noise in the condensational growth equation parameterizes the influence of supersaturation fluctuations and gives analytical formulas for computing DSDs, here identified as Gibbs states; see Eq.
 (9).
-Such probability distribution is a general formulation of DSDs found in the literature McGraw and Liu (2006); Chandrakar et al.
-(2020); Krueger (2020), with the addition of size dependent fluctuations.
-The resulting DSDs are classified into three types according to the mean supersaturation parameter $\lambda$.
-The first corresponds to the case where supersaturation fluctuations do not have the strength to activate droplets for a prolonged period of time and, hence, the size distribution is unimodal around the haze state.
-In the second case, activation is allowed and if the threshold $\lambda_{h}$ is surpassed, haze particles cease to be stable and a single DSD mode is observed peaking over the activated sizes.
-Consequently, even if Köhler’s threshold supersaturation is not attained, there is a finite probability that particles activate and remain so stably.
-The associated timescales for activation is therefore related to the average time it takes to droplets to gain sufficient energy to surpass the potential barrier of activation— see Fig.
-1c)—, and they are given precisely by Kramers’ formula Eq.
-(12).
-In fact, when a cloud parcel is adiabatically cooled, supersaturation fluctuates around a slowly increasing mean.
-Neglecting noisy fluctuations would incur into an overestimation of the superstaruation threshold for activation, as shown in every panel of Fig.
-5, where the interplay between noise and activation hysteresis is illustrated.
-It is realized that on an activation-deactivation cycle, on average, activation occurs at a much lower supersaturations compared to $\lambda_{h}$, and the larger the variance the lower such threshold is.
-The confrontation of our population Köhler scheme within our Brownian modeling framework to steady-state cloud experiments conducted in the Pi-chamber has revealed a striking consistency.
-Particularly at capturing to a high degree of fidelity the complexity of the DSDs’ shapes along with their bimodal features as reported in the experimental study of Prabhakaran et al.
-(2020), for different saturation regimes.
-Under such a setting, it remains to investigate the slow adiabatic changes in mean supersaturation.
-Based on the analytical insights gained on the droplet activation-deactivation hysteresis (see Fig.
-5), we foresee that the same hysteresis behavior should also be observed in the experimental framework of Prabhakaran et al.
-(2020).
-Weak clouds in nature are likely to be exposed to aerosols of several sizes and compositions.
-However, the weak supersaturation may act as a selective filter, making the largest haze bins the leading players, yielding conditions that are not so different from our theoretical setting.
-Moreover, the additional Brownian noise blurs the particles’ homogeneity due to the monodisperse nature of the initial conditions.
-Nevertheless, cases with polydisperse particles of different chemical natures and, hence, different Köhler curves are planned to be investigated.
-METHODS
-Multistability.
-In order for model (3) to allow for multiple stable equilibria, it is enough if the the M-Köhler curve $f-g$ has at least a local minimum and a local maximum.
-This property is to guarantee that the zeros of $\lambda-(f-g)$ correspond to at least two stable equilibria and an unstable one, for a range of values of $\lambda$.
+Such probability distributions generalize earlier formulas for DSDs found in the literature such as Weibull distributions [ML06, CSY+20, Kru20].
+The resulting Droplet Size Distributions (DSDs) are categorized into three types based on the mean supersaturation parameter $\lambda$.
+In the first type, fluctuations in supersaturation are insufficient to sustain droplet activation for extended periods, leading to a unimodal DSD centered around the haze state.
+In the second type, activation is possible, and when the threshold $\lambda_{h}$ is exceeded, haze particles become unstable.
+This results in a single DSD mode peaking at activated sizes.
+Notably, even if Köhler’s threshold supersaturation is not reached, there is a finite probability of particle activation.
+The associated activation timescales are linked to the average time required for droplets to overcome the activation energy barrier (Fig.
+1c), as accurately described by Kramers’ formula (Eq.
+(12)).
+When a cloud parcel undergoes adiabatic cooling, supersaturation fluctuates around a gradually increasing mean.
+Ignoring turbulent fluctuations can lead to an overestimation of the activation threshold, as illustrated in Figure 5.
+This figure depicts various hysteresis loops where activation occurs at even lower supersaturation values than $\lambda_{h}$, especially when the noise variance is large.
+Our stochastic M-Köhler model has demonstrated remarkable agreement with steady-state cloud experiments conducted in the Pi-chamber.
+Notably, it accurately captures the intricate DSD shapes, including bimodal features, as observed in the experimental study by [PSK+20] across various saturation regimes.
+To further explore this consistency, future investigations should focus on slow adiabatic changes in mean supersaturation.
+Based on the analytical understanding of droplet activation-deactivation hysteresis (Fig.
+5), we anticipate that similar hysteresis behavior will be observed in the experimental setup of [PSK+20].
+In natural environments, weak clouds often encounter aerosols of varying sizes and compositions.
+Nevertheless, the low supersaturation levels may act as a selective filter, favoring the activation of larger haze particles.
+This can create conditions that are not substantially different from our theoretical framework.
+Additionally, the presence of Brownian motion can introduce variability into the particle distribution, even with initially monodisperse conditions.
+Future studies will explore scenarios involving polydisperse particles with diverse chemical properties and varying Köhler curves.
+Acknowledgments
+The insights discussed in this work have substantially benefited from the reviewers’ constructive comments, and we express our deep gratitude to them.
+Funding:
+This work has been partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no.
+810370).
+MSG is grateful to the Feinberg Graduate School for their support through the Dean of Faculty Fellowship.
+Author contributions:
+MDC conceived the presented idea.
+MSG led the analyses and MDC supported.
+MSG, MDC, and IK discussed the results and wrote the manuscript.
+All co-authors provided critical feedback and helped shape the research, analysis and manuscript.
+Competing interests:
+The authors declare no competing interests.
+Data and materials availability:
+All experimental data used in this study are based on the article of [PSK+20] and are publicly available at: https://digitalcommons.mtu.edu/data-files/3/.
+All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
+Material and Methods
+Multistability
+To ensure that the model defined by Equation (3) exhibits multiple stable equilibria, it suffices for the M-Köhler curve $f-g$ to possess at least one local minimum and one local maximum.
+This condition guarantees that the roots of the equation $\lambda-(f-g)=0$ correspond to at least two stable equilibria and one unstable equilibrium within a certain range of $\lambda$ values.
 To simplify the presentation of the results stated in Proposition 1 below, we introduce the following parameters:
@@
 $\tilde{B}$ $=\frac{B}{(2D)^{3/2}}.$ (17b)
-We then have the following sufficiency result for the existence of a local maximum and a local minimum:
+We then have the following sufficiency conditions for the existence of a local maximum and a local minimum:
 Proposition 1.
@@
 (3) has at least a local maximum and a local minimum.
-Proof.
-We need to show that the function $\dot{f}-\dot{g}$ has two zeros, and thus it suffices to prove that $\dot{f}-\dot{g}$ has at least two changes of sign. This function reads, for $X>0$, as:
+Proof. To demonstrate the existence of two zeros for the function $\dot{f}-\dot{g}$, it is sufficient to prove that it undergoes at least two sign changes. For $X>0$, this function is given by:
 $\dot{f}(X)-\dot{g}(X)=-\frac{\tilde{A}}{2}X^{-3/2}+\frac{3\tilde{B}}{2}X^{-5/2}+\alpha\beta X^{\alpha-1}.$ (18)
 When $\alpha=3/2$ as in Eq. (6), one recovers that finding the zeros of $\dot{f}-\dot{g}$ is equivalent to finding the roots of the polynomial equation (7).
-First of all, $\dot{f}$ has a single local and global minimum. Indeed, we see this by examining $\ddot{f}$:
+First of all, $\dot{f}$ has a single local and global minimum. Indeed, this can be observed by examining $\ddot{f}$:
 $\ddot{f}(X)=\frac{3}{4}\tilde{A}X^{-5/2}-\frac{15}{4}\tilde{B}X^{-7/2},$ (19)
@@
 Considering this limit and the one in Eq. (21b), there exists $X_{2}>5\tilde{B}/\tilde{A}$ such that $\dot{f}(X_{2})-\dot{g}(X_{2})>0$. This gives the second change of sign.
-Due to the continuity and smoothness of the functions $f$ and $g$ on $\mathrm{R}^{+}$, the function $\dot{f}-\dot{g}$ must have at least two zeros. ∎
-The confining potential.
+Due to the continuity and smoothness of the functions $f$ and $g$ on $\mathrm{R}^{+}$, the function $\dot{f}-\dot{g}$ must have at least two zeros.
+The confining potential
 In order for a gradient system to possess a Gibbs state, it is necessary that the associated potential is confining see e.g.
-(Pavliotis, 2014, §4).
+Chapter 4 in [Pav14].
 We provide here sufficient mathematical conditions by which a sink function $g$ in Eq.
@@
 (8) is confining.
-Proof.
-By the definition of $f$, $g$ and the potential $V_{\lambda}$, we find that:
+Proof. By the definition of $f$, $g$ and the potential $V_{\lambda}$, we find that:
 $-V$ (${}_{\lambda}(X)=\int^{X}\left(\lambda-f(X)+g(X)\right)\,\mathrm{d}X$ (26a)
@@
 $\frac{\xi\beta}{1+\alpha}X_{0}^{\alpha}-\xi\lambda=1.$ (29)
-This proves the confining property. ∎
-Hysteresis path algorithm.
+This proves the confining property.
+Hysteresis path algorithm
 To produce the hysteresis paths shown in Fig.
@@
 To form the lower branch of this hysteresis path, we start with $\lambda=\lambda_{0}$ and from the haze state $X_{h}(\lambda_{0})$, and solve Eq.
-(6) using an Euler-Maruyama scheme (Kloeden and Platen, , Chapter 9.1) for one timestep; $\mathrm{d}t=1$.
+(6) using an Euler-Maruyama scheme with time-step $\delta t=10^{-2}\mathrm{s}$; see Chapter 9.1 in [KP92].
 Then $\lambda$ is updated to $\lambda_{0}+j\delta\lambda$ and Eq.
@@
 Consequently, the larger $N$ is, the smaller $\delta\lambda$, and vice versa.
-Figure 7: Noise-induced metastable states.
+Figure 7.
+Noise-induced metastable states.
 Panels a), b) and c), correspond to the three different supersaturation configurations as indicated in the titles and corresponding to those of Fig.
@@
 (34) which also correspond to the haze, unstable and activated metastable states, denoted by $d_{h},d_{u}$ and $d_{c}$, respectively.
-The Lamperti transformation.
+The Lamperti transformation
 For scalar stochastic differential model, it is possible to map a stochastic differential equation (SDE) driven by state-dependent noise onto an SDE driven by an additive noise.
@@
 $\widehat{V}(Y)=\frac{V(h^{-1}(Y))}{\sigma(h^{-1}(Y))}-\frac{1}{2}\sigma^{\prime}(h^{-1}(Y)).$ (32)
-This is called the Lamperti transformation (Pavliotis, 2014, §3.6).
+This is called the Lamperti transformation; see Chapter 3.6 in [Pav14].
 The Lamperti transformation allows for identifying the effective potential, $\widehat{V}$, for the stochastic model (30) subject to state-dependent stochastic disturbances given by $\sigma(X)\dot{W}_{t}$.
@@
 We talk then of noise-induced metastability in case of creation of an extra metastable state resulting from the interaction of the drift term with the diffusion term $\sigma(X)\dot{W}_{t}$.
-Noise-induced metastability. By applying (32) to our stochastic droplet model (15) in Pi-chamber, we obtain:
+Noise-induced metastability
+By applying (32) to our stochastic droplet model (15) in Pi-chamber, we obtain:
 $\widehat{V}^{\prime}_{\lambda}(Y)=-\frac{V^{\prime}_{\lambda}(X)}{\sigma(X)}-\frac{1}{2}\sigma^{\prime}(X),\mbox{ with }Y=\int^{X}\sigma^{-1}(X)\,\mathrm{d}X,$ (33)
@@
 (11).
-Thus, if one wishes to match the locations of these metastable states with let us say the modes of an empirical distribution (as in PI-Chamber Empirical Distributions vs Gibbs States), one needs to solve Eq.
+Thus, if one wishes to match the locations of these metastable states with let us say the modes of an empirical distribution (as in Section "Pi-Chamber empirical distributions vs Gibbs states"), one needs to solve Eq.
 (34).
@@
 7 correspond to the left-hand-side (LHS) and right-hand-side (RHS) of Eq.
-(34), respectively, for the three case studies of PI-Chamber Empirical Distributions vs Gibbs States.
-The intersection of these two curves determines the location of the metastable states indicated by the dashed vertical lines for Cases I, II and III with the parameter configuration shown in Table 2.
-Acknowledgements This work has been partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no.
-810370).
-MSG is grateful to the Feinberg Graduate School for their support through the Dean of Faculty Fellowship.
-Author contributions: MDC conceived the presented idea.
-MSG led the analyses and MDC supported.
-MSG, MDC, OA, and IK discussed the results and wrote the manuscript.
-All co-authors provided critical feedback and helped shape the research, analysis and manuscript.
-Competing interests: The authors declare no competing interests.
-Data availability: All experimental data used in this study are based on the article of Prabhakaran et al.
-(2020) and are publicly available at: https://digitalcommons.mtu.edu/data-files/3/.
-The numerical data obtained for this article are available upon request to the authors.
+(34), respectively, for the three case studies of Section "Pi-Chamber empirical distributions vs Gibbs states." The intersection of these two curves determines the location of the metastable states indicated by the dashed vertical lines for Cases I, II and III with the parameter configuration shown in Table 2.